### v1

# On lower bounds for the distances between APN functions

Abstract
Whether two distinct APN functions can have a Hamming distance of $1$ remains an open problem. In 2020, L. Budaghyan et al. introduced a new CCZ-invariant $\Pi_{F}$ which can be used to provide lower bounds on the Hamming distance between a given APN function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ and other APN functions. Lower bounds on the distance from an APN function $F$ to any other are known for almost bent (AB) functions and when $F$ is a 3-to-1 quadratic function with $n$ even. In this paper, we reinterpret $\Pi_{F}$ in terms of the exclude multiplicities of the graph $\mathcal{G}_{F}=\left\{(x,F(x)):x\in\mathbb{F}_{2}^{n}\right\}$ of $F$ as a Sidon set. We establish lower bounds on the distance for even $n$ when $F$ is plateaued APN, generalize the known lower bounds for quadratic 33-to-11 function to all 3-to-1 plateaued functions (e.g. Kasami functions), and derive new lower bounds for when $F$ is the APN inverse function over $\mathbb{F}_{2^{n}}$ for $n$ odd. We also study how the exclude multiplicities of $\mathcal{G}_{F}$ are directly connected to the existence of linear structures of $\gamma_{F}$ when $F$ is plateaued and APN and the ortho-derivative when $F$ is a quadratic APN function. We also use the CCZ-invariance of exclude multiplicities to prove that the Brinkmann-Leander-Edel-Pott function is not CCZ-equivalent to a plateaued function.
Keywords Almost perfect nonlinear (APN) functions, Sidon sets, vectorial Boolean functions, Hamming distance
Mathematics Subject Classification (2020) Primary: 94D10, 11B13.
Secondary: 94A60.
1 Introduction
We call a function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{m}$ a vectorial Boolean function or an $(n,m)$-function. Vectorial functions are prevalent and one of the main focuses of study in cryptography, and in particular, they are used as the primary nonlinear components of block ciphers. In this paper, we are concerned with the $(n,n)$-functions that are optimally resistant to a differential attack, an attack on a block cipher first introduced by Biham and Shamir in [2]. Such functions are called almost perfect nonlinear (APN). In particular, we call $F$ APN if $F(x)+F(x+a)=b$ has either 0 or $2$ solutions for all $a,b\in\mathbb{F}_{2}^{n}$ where $a\neq 0$.
An active area of research is the construction of new APN functions, and in this paper, we are concerned with obtaining APN functions from already-known APN functions. In particular, it is conjectured that changing the value of an APN function at a single point results in a function that is no longer APN [6]. For two vectorial Boolean functions $F,G\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$, we define the Hamming distance between $F$ and $G$ as
$d(F,G)=|\left\{x\in\mathbb{F}_{2}^{n}:F(x)\neq G(x)\right\}|.$
Said differently, it is conjectured that any two distinct APN functions have Hamming distance greater than $1$, and it is known that this conjecture holds in certain cases, such as when $F$ is a plateaued APN function (see Section 2). Moreover, if two APN functions have a distance of $1$, at least one of them has algebraic degree $n$ (see [12]), and it is also conjectured that no APN function has algebraic degree $n$ [6].
In [5], the CCZ-invariant $\Pi_{F}$ was first introduced. It was shown that $\Pi_{F}$ can be used to determine a lower bound on the Hamming distance between distinct APN functions $F$ and $G$. From this, the authors of [5] derived lower bounds on the distance to the Gold function $F(x)=x^{3}$, and they observed that their lower bound tends to infinity as $n$ grows [5]. In the case that $n$ is odd, the same bounds for $F(x)=x^{3}$ of [5] were generalized to all AB functions in [17]. Moreover, in the case that $n$ is even, the same bounds for $F(x)=x^{3}$ of [5] were generalized to all quadratic 33-to-11 functions in [7]. Besides AB functions and quadratic 3-to-1 functions, no other strong lower bounds between distances of APN functions have been known until this paper.
In this paper, we have three main results. The first is a new lower bound on distance for plateaued APN functions which significantly improves on the previous lower bound of $d(F,G)\geq 2$. We then generalize the result of [7] from quadratic 3-to-1 functions to all plateaued 3-to-1 functions (which are necessarily APN). The third main result is regarding the APN inverse function, and we provide the first known lower bounds in terms of binary Kloosterman sums $K_{n}(a)=\sum_{x\in\mathbb{F}_{2}^{n}}(-1)^{\operatorname{Tr}(ax+x^{-1})}$.
Theorem 1.1.
Suppose $n\geq 4$ is even, let $F,G\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be distinct APN functions such that $F$ is plateaued. Then $d(F,G)\geq 2^{\frac{n}{2}-1}$.
Theorem 1.2.
Suppose $n\geq 4$ is even, let $F,G\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be distinct APN functions such that $F$ is plateaued and 33-to-11. Then
$d(F,G)\geq\begin{cases}\frac{1}{3}(2^{n-1}-2^{\frac{n}{2}}+2)&n\equiv 0\mod 4,\\ \frac{1}{3}(2^{n-1}-2^{\frac{n}{2}-1}+2)&n\equiv 2\mod 4.\end{cases}$
Theorem 1.3.
Suppose $n\geq 3$ is odd. Let $F\colon\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ be the function $F(x)=x^{-1}$, where we define $\frac{1}{0}:=0$. If $G\colon\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ is an APN function such that $G\neq F$, then
$d(F,G)\geq\left\lceil\frac{2^{n}-5-\max_{a\in\mathbb{F}_{2^{n}}^{\ast}}|K_{n}(a)-1|}{6}\right\rceil+1\approx\frac{2^{n}-2^{\frac{n}{2}+1}}{6}+1.$
Our proofs of the above results are all done via a reinterpretation of the lower bound in [5] in the language of (binary) Sidon sets and their so-called exclude multiplicities. A Sidon set in $\mathbb{F}_{2}^{n}$ is a subset $S\subseteq\mathbb{F}_{2}^{n}$ such that no four distinct points in $S$ have zero sum. APN functions and Sidon sets are closely related as the graph $\mathcal{G}_{F}=\left\{(x,F(x)):x\in\mathbb{F}_{2}^{n}\right\}$ of a function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is a Sidon set in $(\mathbb{F}_{2}^{n})^{2}$ if and only if $F$ is APN. For a Sidon set $S\subseteq\mathbb{F}_{2}^{n}$ and a point $p\in\mathbb{F}_{2}^{n}\setminus S$, we say that the exclude multiplicity $\operatorname{mult}_{S}(p)$ of $p$ is
$\operatorname{mult}_{S}(p)=|\left\{\left\{x,y,z\right\}\subseteq S:x+y+z=p\right\}|.$
If $\operatorname{mult}_{S}(p)>0$, we call $p$ an exclude point of $S$. We call $S$ a maximal Sidon set if $S$ is a Sidon set such that any point in $\mathbb{F}_{2}^{n}\setminus S$ has nonzero exclude multiplicity as such a set cannot be contained in a strictly larger Sidon set. Carlet demonstrated in [12] that the graph of any APN function is a maximal Sidon set if and only if any two distinct APN functions have a Hamming distance strictly greater than $1$.
For an APN function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$, we will observe that the $\Pi_{F}$ CCZ-invariant from [5] has a very natural connection to the exclude multiplicities of $\mathcal{G}_{F}$. More precisely, if $F$ is APN, then we have the following relation:
$\Pi_{F}=\left\{2^{n}\right\}\cup\left\{3\operatorname{mult}_{\mathcal{G}_{F}}(a,b):(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}\right\}.$
In [5], it was shown that if $F$ is APN, then $d(F,G)\geq\lceil\frac{\min\Pi_{F}}{3}\rceil+1$, where $G\neq F$ is APN. Hence, for an APN function $F$, this inequality can then be expressed in terms of exclude multiplicity:
$d(F,G)\geq e_{\min}(\mathcal{G}_{F})+1,$
where $e_{\min}(\mathcal{G}_{F})=\min_{(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}}\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$. The proofs of each of our main results all rely on finding a lower bound on $e_{\min}(\mathcal{G}_{F})$, and in some cases, we determine the exact exclude multiplicities of $\mathcal{G}_{F}$.
$F\colon\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ Lower bounds of $d(F,G)$ for $G\neq F$ and $G$ APN Reference
$F$ almost bent (AB) $\frac{1}{3}(2^{n-1}+2)$ [17], Theorem 4.2
$F$ plateaued APN, $n\geq 4$ even $2^{\frac{n}{2}-1}$ Theorem 1.1
$F$ 3-to-1 plateaued, $n\geq 4$ even $\begin{cases}\frac{1}{3}(2^{n-1}-2^{\frac{n}{2}}+2)&n\equiv 0\mod 4,\\ \frac{1}{3}(2^{n-1}-2^{\frac{n-2}{2}}+2)&n\equiv 2\mod 4.\end{cases}$ Theorem 1.2
$F(x)=x^{-1}$, $n\geq 3$ odd $\left\lceil\frac{1}{6}\left(2^{n}-5-\max_{a\in\mathbb{F}_{2^{n}}^{\ast}}|K_{n}(a)-1|\right)\right\rceil+1$ Theorem 1.3
Table 1: For an APN function $F$ in a given family, its distance is greater than or equal to the value in the second column to all other APN functions $G\neq F$.
The remainder of this paper is organized as follows. In Section 2, we provide the necessary preliminary and background for this paper. In Section 3, we recall results of [5] and express $\Pi_{F}$ in terms of exclude multiplicity.
In Section 4, we study the exclude multiplicities of the graphs of plateaued APN functions, and we prove Theorems 1.1 and 1.2. Since all plateaued APN functions are AB when $n$ is odd, we mostly focus on the case when $n$ is even. We express the exclude multiplicities of $\mathcal{G}_{F}$ in terms of a sum involving the amplitudes of the components of $F$. Moreover, we generalize a result of [5] to all plateaued functions, showing that $\Pi_{F}$ is completely determined by $\Pi_{F}^{0}$ (see Section 3) when $F$ is plateaued and APN. As a corollary, we prove that the Brinkmann-Leander-Edel-Pott function (the only known instance of an APN function not CCZ-equivalent to a power function or a quadratic function) is also not CCZ-equivalent to a plateaued function. In Section 4.4, we study quadratic APN functions with a focus on how their ortho-derivatives can be used to express exclude multiplicities. By using a recently proved fact about the algebraic degree of an ortho-derivative, we show that all exclude multiplicities of $\mathcal{G}_{F}$ are odd when $F$ is quadratic APN. As an equivalent result, we prove that ${\mathcal{B}({F})}=\left\{v\in\mathbb{F}_{2}^{n}:v\cdot F\text{ is bent}\right\}$ always has an intersection of even size with any affine hyperplane of $\mathbb{F}_{2}^{n}$ when $F$ is quadratic APN.
In Section 5, we study the exclude multiplicities of APN power functions, and we show that $\Pi_{F}$ is completely determined by $\Pi_{F}^{0}$ and $\Pi_{F}^{1}$. In particular, we show that $\Pi_{F}^{\beta}$ does not depend on $\beta\in\mathbb{F}_{2^{n}}^{\ast}$. Then, in Section 6, we prove Theorem 1.3. In particular, in Section 6, we use results of Lachaud and Wolfmann in [32] on elliptic curves over $\mathbb{F}_{2^{n}}$ and their connection to binary Kloosterman sums in order to derive our lower bounds in the case of $F(x)=x^{-1}$ when $n$ is odd.
In Section 7, we study how exclude multiplicity is related to the $\gamma_{F}$ Boolean function associated to an APN function $F$. In particular, we show that if $F$ is plateaued APN, then $\gamma_{F}$ has no nontrivial linear structures if and only if $\mathcal{G}_{F}$ does not have a point of exclude multiplicity $\frac{2^{n}-1}{3}$ (the largest possible). As a consequence, we prove that $\gamma_{F}$ has no nontrivial linear structures when $F$ is plateaued 3-to-1. We also show that $\gamma_{F}$ has no nontrivial linear structures when $F$ is quadratic APN with an ortho-derivative of nonzero nonlinearity. To conclude the paper, we list open problems in Section 8.
2 Preliminaries
For a vectorial Boolean function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$, we call $D_{a}F(x)=F(x)+F(x+a)$ the derivative of $F$ in the direction of $a$. We say that $F$ is almost perfect nonlinear (APN) if $D_{a}F(x)=b$ has at most $2$ solutions for all $a\in\mathbb{F}_{2}^{n}\setminus\left\{0\right\}$ and all $b\in\mathbb{F}_{2}^{n}$. APN functions are of interest in cryptography as they provide optimal resistance to a differential attack (first introduced in [2]) when used as a substitution box (S-box) in a block cipher [33].
For a vectorial Boolean function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$, the Walsh transform of $F$ is defined as $W_{F}(u,v)=\sum_{x\in\mathbb{F}_{2}^{n}}(-1)^{x\cdot u+F(x)+v}$, where $\cdot$ is the standard inner product on $\mathbb{F}_{2}^{n}$. It is a classical result of [16] that $F$ is APN if and only if $\sum_{(u,v)\in(\mathbb{F}_{2}^{n})^{2}}W_{F}^{4}(u,v)=2^{2n}(3\cdot 2^{2n}-2^{n+1})$. We say that the linearity of $F$ is the value $\mathcal{L}(F)=\max_{(u,v)\in(\mathbb{F}_{2}^{n})^{2}\setminus(0,0)}|W_{F}(u,v)|$, and the nonlinearity of $F$ is $\operatorname{\mathcal{NL}}(F)=2^{n-1}-\frac{1}{2}\mathcal{L}(F)$. Oftentimes, we identify $\mathbb{F}_{2}^{n}$ with $\mathbb{F}_{2^{n}}$, and in which case, we define $u\cdot v$ as $\operatorname{Tr}(uv)$ where $\operatorname{Tr}$ is the absolute trace function $\operatorname{Tr}(x)=\sum_{i=0}^{n-1}x^{2^{i}}$ from $\mathbb{F}_{2^{n}}$ onto $\mathbb{F}_{2}$. The Walsh transform of $F$ can also be expressed as the Fourier-Hadamard transform of the indicator $1_{\mathcal{G}_{F}}$ of $\mathcal{G}_{F}$. Recall that the Fourier-Hadamard transform of a function $\varphi\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}$ is the function $\widehat{\varphi}(u)=\sum_{x\in\mathbb{F}_{2}^{n}}(-1)^{x\cdot u}\varphi(x)$. We will also use the convolutional product, which for any two functions $\varphi,\psi\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}$ is given by $(\varphi\times\psi)(u)=\sum_{x\in\mathbb{F}_{2}^{n}}\varphi(x)\psi(x+u)$.
For a Boolean function $f\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$, we say that its (Hamming) weight is the size of its support, i.e. $\operatorname{wt}(f)=\sum_{x\in\mathbb{F}_{2}^{n}}f(x)$. We say that $f$ is balanced if $\operatorname{wt}(f)=2^{n-1}$. We also define the linearity $\mathcal{L}(f)$ of $f$ as the maximum absolute value of its Walsh transform $W_{f}(u)=\sum_{x\in\mathbb{F}_{2}^{n}}(-1)^{f(x)+x\cdot u}$, and the nonlinearity of $f$ is $\operatorname{\mathcal{NL}}(f)=2^{n-1}-\frac{1}{2}\mathcal{L}(f)$. A function is bent if and only if its linearity is minimal with $\mathcal{L}(f)=2^{\frac{n}{2}}$, and so bent functions only exist when $n$ is even. Equivalently, $f$ is bent if and only if $D_{a}f$ is balanced for all nonzero $a\in\mathbb{F}_{2}^{n}$. For a given vectorial Boolean function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ and $v\in\mathbb{F}_{2}^{n}$, we say that $(v\cdot F)(x)=v\cdot F(x)$ is a component function of $F$, and we denote the set of all bent component functions of $F$ as $\mathcal{B}({F})$}.
Another important class of vectorial Boolean functions are those such that $W_{F}(u,v)\in\left\{0,\pm 2^{\frac{n+1}{2}}\right\}$ for all $(u,v)\in(\mathbb{F}_{2}^{n})^{2}\setminus\left\{(0,0)\right\}$. If $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ has this property, then we call $F$ almost bent (AB). Note that AB functions only exist when $n$ is odd as the Walsh transform always takes integer values, and any AB function is APN (see [16] or [35], for instance).
We call two vectorial Boolean functions functions $F,G\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ CCZ-equivalent if there exists an affine permutation $A\colon(\mathbb{F}_{2}^{n})^{2}\to(\mathbb{F}_{2}^{n})^{2}$ such that $A(\mathcal{G}_{F})=\mathcal{G}_{G}$. It is conjectured that all APN functions of the form $F(x)=x^{d}$, defined over $\mathbb{F}_{2^{n}}$, are known up to CCZ-equivalence [22]. In Table 2 and Table 3, we list all known families of APN and AB monomials over $\mathbb{F}_{2^{n}}$.
Name $d$ Condition Reference(s)
Gold $2^{k}+1$ $\gcd(k,n)=1$ [24, 33]
Kasami $2^{2k}-2^{k}+1$ $\gcd(k,n)=1$ [26, 28]
Welch $2^{t}+3$ $n=2t+1$ [20]
Niho $\begin{cases}2^{t}+2^{\frac{t}{2}}-1&\text{if $t$ even}\\ 2^{t}+2^{\frac{3t+1}{2}}-1&\text{if $t$ odd}\end{cases}$ $n=2t+1$ [22]
Inverse $2^{2t}-1$ $n=2t+1$ [33, 1]
Dobbertin $2^{4t}+2^{3t}+2^{2t}+2^{t}-1$ $n=5t$ [21]
Table 2: Known infinite families of APN power functions $\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ of the form $x\mapsto x^{d}$.
Name $d$ Condition Reference(s)
Gold $2^{k}+1$ $\gcd(k,n)=1$ [24, 33]
Kasami $2^{2k}-2^{k}+1$ $\gcd(k,n)=1$ [28]
Welch $2^{t}+3$ $n=2t+1$ [9, 8]
Niho $\begin{cases}2^{t}+2^{\frac{t}{2}}-1&\text{if $t$ even}\\ 2^{t}+2^{\frac{3t+1}{2}}-1&\text{if $t$ odd}\end{cases}$ $n=2t+1$ [25]
Table 3: Known infinite families of AB power functions $\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ of the form $x\mapsto x^{d}$, $n$ odd.
Also of importance are plateaued functions, which are not necessarily APN. We call a function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ plateaued if for all $v\in\mathbb{F}_{2}^{n}$, there exists a positive integer $\lambda_{v}\geq 0$ such that $W_{F}(u,v)\in\left\{0,\pm\lambda_{v}\right\}$ for all $u\in\mathbb{F}_{2}^{n}$. For example, any AB function is plateaued by definition. APN plateaued functions and APN power functions comprise almost all known families of known APN functions.
It is unknown if there exists two APN functions $F,G\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ of Hamming distance $1$ from each other. In particular, we define the Hamming distance $d(F,G)$ as
$d(F,G)=|\left\{x\in\mathbb{F}_{2}^{n}:F(x)\neq G(x)\right\}|.$
There is the following conjecture on the distance between APN functions.
Conjecture 2.1 ([6]).
Assume $n\geq 3$. Suppose $F,G\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ are APN functions such that $F\neq G$. Then $d(F,G)>1$.
In general, not much progress has been made towards this conjecture for all APN functions, but it was shown to be true in the cases when $F$ is a known APN power function (besides the Dobbertin function) or APN plateaued [6].
APN functions also have a natural connection to Sidon sets as the graph of $F$ is a Sidon set in $(\mathbb{F}_{2}^{n})^{2}$ if and only if $F$ is APN. A Sidon set is a set $S\subseteq\mathbb{F}_{2}^{n}$ such that $x+y+z+w=0$ has no solutions for all pairwise distinct $x,y,z,w\in S$. A Sidon set is called maximal if it is not contained in any strictly larger Sidon set. In [12] used the connection of APN functions and Sidon sets to prove that 2.1 is equivalent to the following conjecture.
Conjecture 2.2 ([12]).
Assume $n\geq 3$, and let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be an APN function. Then $\mathcal{G}_{F}$ is a maximal Sidon set.
Sidon sets that are maximal can also be characterized by their exclude multiplicities. If $S$ is a Sidon set and $p\in\mathbb{F}_{2}^{n}\setminus S$, then we say that the exclude multiplicity $\operatorname{mult}_{S}(p)$ of $p$ (with respect to $S$) is the number of distinct triples contained in $S$ that sum to $p$, i.e.
$\operatorname{mult}_{S}(p)=|\left\{\left\{x,y,z\right\}\subseteq S:x+y+z=p\right\}|.$
We call a point $p\notin S$ of nonzero exclude multiplicity an exclude point of $S$. Hence, a Sidon set $S\subseteq\mathbb{F}_{2}^{n}$ is maximal if and only if every point in $\mathbb{F}_{2}^{n}\setminus S$ has nonzero exclude multiplicity with respect to $S$ (cf. [19]). In other words, if $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is an APN function, then it is conjectured that $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\neq 0$ for all $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$.
For a set $S\subseteq\mathbb{F}_{2}^{n}$, define $\delta_{S}\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}_{\geq 0}$ to be the function given by $\delta_{S}(a)=|S\cap(a+S)|$ for all $a\in\mathbb{F}_{2}^{n}$. Also, define $\gamma_{S}\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ as the Boolean function
$\gamma_{S}(a)=\begin{cases}1&a\neq 0\text{ and }\delta_{S}(a)>0,\\ 0&\text{otherwise}.\end{cases}$
When $n$ is even and $S=\mathcal{G}_{F}$ for some vectorial Boolean function $F\colon\mathbb{F}_{2}^{n/2}\to\mathbb{F}_{2}^{n/2}$, we simply denote $\delta_{S}$ as $\delta_{F}$ and $\gamma_{S}$ as $\gamma_{F}$. A set $S\subseteq\mathbb{F}_{2}^{n}$ is Sidon if and only if $\delta_{S}(a)\leq 2$ for all nonzero $a\in\mathbb{F}_{2}^{n}$.
3 Distance between APN functions
In this section, we establish the following lemma in two ways: first we prove it using the notion of exclude multiplicity, and then we demonstrate that the notation and methods of [5] can be also used to derive the same bound.
Lemma 3.1.
Let $F,G\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be APN functions such that $F\neq G$. Then
$d(F,G)\geq e_{\min}(\mathcal{G}_{F})+1,$
where $e_{\min}(\mathcal{G}_{F})$ denotes $\min_{(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}}\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$.
Proof.
Suppose $F,G\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ are distinct APN functions. Let $x\in\mathbb{F}_{2}^{n}$ such that $F(x)\neq G(x)$, and let $k=\operatorname{mult}_{\mathcal{G}_{F}}(x,G(x))$. If $k=0$, then $d(F,G)\geq 1=e_{\min}(\mathcal{G}_{F})+1$ by assumption. So, assume $k>0$. By definition, there exist distinct (and necessarily pairwise disjoint) triples $\left\{(x_{i},F(x_{i})),(y_{i},F(y_{i})),(z_{i},F(z_{i}))\right\}_{i=1}^{k}$ such that $(x_{i}+y_{i}+z_{i},F(x_{i})+F(y_{i})+F(z_{i}))=(x,G(x))$ for all $1\leq i\leq k$. For any $1\leq i\leq k$, the equation $x_{i}+y_{i}+z_{i}=x$ implies that $x\notin\left\{x_{i},y_{i},z_{i}\right\}$ as $|\left\{x_{i},y_{i},z_{i}\right\}|=3$. Also, $\mathcal{G}_{G}$ is Sidon, so $\left\{(x_{i},F(x_{i})),(y_{i},F(y_{i})),(z_{i},F(z_{i}))\right\}$ is not contained in $\mathcal{G}_{G}$ for any $1\leq i\leq k$. Hence, there exist $a_{1},\dots,a_{k}$ such that $a_{i}\in\left\{x_{i},y_{i},z_{i}\right\}$ for all $1\leq i\leq k$ and $(a_{i},F(a_{i}))\notin\mathcal{G}_{G}$. Therefore, $F$ and $G$ differ at $x,a_{1},\dots,a_{k}$, implying $d(F,G)\geq k+1\geq e_{\min}(\mathcal{G}_{F})+1$. ∎
Remark 3.2.
Note that Lemma 3.1 can be generalized to Sidon sets with the same bound except the +1+1 term. Let $S,T\subseteq\mathbb{F}_{2}^{n}$ be distinct Sidon sets. Let $y\in T$ such that $y\notin S$, and let $k=\operatorname{mult}_{S}(y)$. By definition, there exist distinct (and necessarily pairwise disjoint) triples $\left\{a_{i},b_{i},c_{i}\right\}_{i=1}^{k}$ such that $a_{i},b_{i},c_{i}\in S$ and $a_{i}+b_{i}+c_{i}=y$. Note that $\left\{a_{i},b_{i},c_{i}\right\}$ cannot be a subset of $T$ for any $1\leq i\leq k$ because this would contradict $T$ being Sidon. Hence, there are at least $k$ points in $S$ that do not lie in $T$, and so $|S\setminus T|\geq k\geq e_{\min}(S)$. Note that we cannot improve this bound to also include the +1+1 term as there exist distinct Sidon sets $S,T\subseteq\mathbb{F}_{2}^{n}$ such that $|S\setminus T|=e_{\min}(S)$.
Let us now recall some notation and definitions that were first introduced in [5]. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a vectorial Boolean function. The shifted derivative $D_{a}^{\beta}F$ of $F$ in direction $a$ with shift $\beta$ is the function
$D_{a}^{\beta}F=D_{a}F(x)+F(a+\beta)=F(x)+F(a+x)+F(a+\beta).$
For an $(n,n)$-function $F$, we denote the image of $D_{a}^{\beta}F$ as $H_{a}^{\beta}F$. Moreover, we denote by $\Pi_{F}^{\beta}(b)$ the set $\Pi_{F}^{\beta}(b)=\left\{a\in\mathbb{F}_{2}^{n}:b\in H_{a}^{\beta}F\right\}$. Also, let
$\Pi_{F}^{\beta}=\left\{|\Pi_{F}^{\beta}(b)|:b\in\mathbb{F}_{2}^{n}\right\},$
and let
$\Pi_{F}=\bigcup_{\beta\in\mathbb{F}_{2}^{n}}\Pi_{F}^{\beta}=\left\{|\Pi_{F}^{\beta}(b)|:\beta,b\in\mathbb{F}_{2}^{n}\right\}.$
We now compute the size of $\Pi_{F}^{\beta}(b)$ in terms of $\gamma_{F}$, and this will be useful in demonstrating the connection between shifted derivatives and exclude multiplicity. Note that for any $u\in\mathbb{F}_{2}^{n}$, we write the Kronecker delta as $\delta_{0}(u)$, which takes value $1$ if and only if $u=0$ and value 0 otherwise.
Proposition 3.3.
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be an APN function. For any $(b,\beta)\in(\mathbb{F}_{2}^{n})^{2}$, we have the equalities
$|\Pi_{F}^{\beta}(b)|$ $=\delta_{0}(b+F(\beta))+\sum_{a\in\mathbb{F}_{2}^{n}}\gamma_{F}(a+\beta,F(a)+b)$
$=\frac{1}{2^{2n+1}}\widehat{W_{F}^{3}}(\beta,b)-\delta_{0}(b+F(\beta))(2^{n-1}-1).$
In particular, if $b\neq F(\beta)$, then $|\Pi_{F}^{\beta}(b)|=3\operatorname{mult}_{\mathcal{G}_{F}}(\beta,b)$.
Proof.
Notice that for any $\beta,b\in\mathbb{F}_{2}^{n}$ , we have
$\Pi_{F}^{\beta}(b)$ $=\left\{a\in\mathbb{F}_{2}^{n}:b\in H_{a}^{\beta}F\right\}$
$=\left\{a\in\mathbb{F}_{2}^{n}:b\in\operatorname{Im}D_{a}F+F(a+\beta)\right\}$
$=\left\{a\in\mathbb{F}_{2}^{n}:\delta_{F}(a,b+F(a+\beta))\neq 0\right\}.$
Therefore the size of $\Pi_{F}^{\beta}(b)$ is then the same as
$\delta_{0}(b+F(\beta))+\sum_{a\in\mathbb{F}_{2}^{n}}\gamma_{F}(a,b+F(a+\beta))=\delta_{0}(b+F(\beta))+\sum_{a\in\mathbb{F}_{2}^{n}}\gamma_{F}(a+\beta,F(a)+b).$
Hence $|\Pi_{F}^{\beta}(b)|=\delta_{0}(b+F(\beta))+\sum_{a\in\mathbb{F}_{2}^{n}}\gamma_{F}(a+\beta,F(a)+b)$. Since $F$ is APN, we know that $\gamma_{F}=\frac{1}{2}\delta_{F}-2^{n}\delta_{(0,0)}$. Also, we have the equality $\delta_{F}=1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}}$ (cf. [16]). Therefore
$\sum_{a\in\mathbb{F}_{2}^{n}}\gamma_{F}(a+\beta,F(a)+b)$ $=(\gamma_{F}\otimes 1_{\mathcal{G}_{F}})(\beta,b)$
$=\frac{1}{2}((\delta_{F}-2^{n}\delta_{0}))\otimes 1_{\mathcal{G}_{F}})(\beta,b)$
$=\frac{1}{2}((1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})\otimes 1_{\mathcal{G}_{F}})(\beta,b)-2^{n-1}1_{\mathcal{G}_{F}}(\beta,b)$
Since $\widehat{\varphi\otimes\psi}=\widehat{\varphi}\widehat{\psi}$ for any functions $\varphi,\psi\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}$ (see [11, Proposition 11]) and using the fact that $\widehat{1_{\mathcal{G}_{F}}}=W_{F}$, the Fourier-Hadamard transform of $(1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})\otimes 1_{\mathcal{G}_{F}}$ is $\widehat{(1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})}\widehat{1_{\mathcal{G}_{F}}}=(\widehat{1_{\mathcal{G}_{F}}}\widehat{1_{\mathcal{G}_{F}}})\widehat{1_{\mathcal{G}_{F}}}=W_{F}^{3}$. Therefore, taking the Fourier-Hadamard transform of both sides, we have $(1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})\otimes 1_{\mathcal{G}_{F}}=\frac{1}{2^{2n}}\widehat{W_{F}^{3}}$, and so
$|\Pi_{F}^{\beta}(b)|$ $=\delta_{0}(b+F(\beta))+\sum_{a\in\mathbb{F}_{2}^{n}}\gamma_{F}(a+\beta,F(a)+b)$
$=\delta_{0}(b+F(\beta))+\frac{1}{2}((1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})\otimes 1_{\mathcal{G}_{F}})(\beta,b)-2^{n-1}1_{\mathcal{G}_{F}}(\beta,b)$
$=\frac{1}{2^{2n+1}}\widehat{W_{F}^{3}}(\beta,b)-\delta_{0}(b+F(\beta))(2^{n-1}-1).$
In particular, when $(\beta,b)\notin\mathcal{G}_{F}$ and $F$ is APN, we have
$|\Pi_{F}^{\beta}(b)|=\frac{1}{2^{2n+1}}\widehat{W_{F}^{3}}(\beta,b)=3\operatorname{mult}_{\mathcal{G}_{F}}(\beta,b).$
Therefore, for an APN function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$, the exclude points of $\mathcal{G}_{F}$ and their multiplicities are fully described by the sizes of the sets $\Pi_{F}^{\beta}(b)$. Note that the set $\Pi_{F}^{\beta}$ describes the exclude multiplicities that occur in $\left\{\beta\right\}\times\mathbb{F}_{2}^{n}$. We have the following corollary of Proposition 3.3.
Corollary 3.4.
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be an APN function. Then
$\Pi_{F}^{\beta}=\left\{2^{n}\right\}\cup\left\{3\operatorname{mult}_{\mathcal{G}_{F}}(\beta,b):b\in\mathbb{F}_{2}^{n}\setminus\left\{F(\beta)\right\}\right\}.$
If $F$ is APN, then the set $\Pi_{F}\setminus\left\{2^{n}\right\}$ is the set of all exclude multiplicities of $\mathcal{G}_{F}$ scaled by $3$. In [5], it was shown that $\Pi_{F}$ can be used to derive a lower bound on the distance from an APN function $F$ to any other APN function.
Corollary 3.5 ([5]).
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be an APN function, and let $m_{F}$ be the number
$m_{F}=\min\Pi_{F}=\min_{b,\beta\in\mathbb{F}_{2}^{n}}|\Pi_{F}^{\beta}(b)|.$
Then for any APN function $G\neq F$ over $\mathbb{F}_{2}^{n}$, the Hamming distance between $F$ and $G$ satisfies
$d(F,G)\geq\left\lceil\frac{m_{F}}{3}\right\rceil+1.$
For any APN function $F$, any exclude point of $\mathcal{G}_{F}$ has exclude multiplicity at most $\lfloor\frac{2^{n}}{3}\rfloor$ (see [19, Corollary 3.4]). Therefore, $m_{F}=\min\left\{3e_{\min}(\mathcal{G}_{F}),2^{n}\right\}=3e_{\min}(\mathcal{G}_{F})$, and so the bound of Corollary 3.5 is equivalent to the bound of Lemma 3.1.
4 Plateaued APN functions
Recall that a Boolean function $f\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ is plateaued if there exists an integer $\lambda\geq 0$ (called the amplitude of $f$) such that $W_{f}(u)\in\left\{0,\pm\lambda\right\}$ for all $u\in\mathbb{F}_{2}^{n}$. A vectorial Boolean function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is called plateaued if all of its components $v\cdot F$ are plateaued, and we denote the amplitude of $v\cdot F$ as $\lambda_{v}$. Almost all of the known APN functions are plateaued because they are almost all quadratic. A function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is quadratic if and only if $D_{a}F(x)=F(x)+F(x+a)$ is affine for all $a\in\mathbb{F}_{2}^{n}$. It is well-known that any quadratic vectorial Boolean function is also plateaued.
To provide a brief overview of this section, we study the exclude multiplicities of $\mathcal{G}_{F}$ when $F$ is a plateaued function and consider particular subclasses of plateaued functions such as AB functions, 3-to-1 plateaued functions, and quadratic APN functions.
4.1 AB functions
The first class of plateaued APN functions that we consider is the class of almost bent (AB) functions. Recall that $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is AB if $W_{F}(u,v)=\left\{0,\pm 2^{\frac{n+1}{2}}\right\}$ for all nonzero $(u,v)\in(\mathbb{F}_{2}^{n})^{2}$, and so AB functions only exist when $n$ is odd. Moreover, all plateaued APN functions are necessarily AB when $n$ is odd. We can easily derive a lower bound on the distance from any AB function to another APN function using the van Dam and Fon-Der-Flaass characterization of AB functions.
Theorem 4.1 ([35]).
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a function. Then $F$ is AB if and only if the system of equations
$\begin{cases}x+y+z=a\\ F(x)+F(y)+F(z)=b\end{cases}$
has $2^{n}-2$ or $3\cdot 2^{n}-2$ solutions $(x,y,z)\in(\mathbb{F}_{2}^{n})^{3}$ for every $(a,b)\in(\mathbb{F}_{2}^{n})^{2}$. If so, then the system has $2^{n}-2$ solutions if $b\neq F(a)$ and $3\cdot 2^{n}-2$ solutions otherwise.
This useful characterization immediately tells us that the exclude multiplicity of any point in $(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$ is $\frac{2^{n}-2}{6}$ when $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is an AB function. The following result immediately follows and was first stated by Coulter and Kaleyski in [17].
Theorem 4.2.
Suppose $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is an AB function, and let $G\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be an APN function such that $G\neq F$. Then
$d(F,G)\geq\frac{2^{n}-2}{6}+1=\frac{2^{n-1}+2}{3}.$
Proof.
Apply Lemma 3.1 to Theorem 4.1. ∎
4.2 Properties of the exclude multiplicities of $\mathcal{G}_{F}$ when $F$ is plateaued APN
To begin, let us first recall the following characterization of plateaued functions.
Theorem 4.3 ([10]).
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{m}$ be a vectorial function. Then
  1. 1.
    $F$ is plateaued if and only if, for every $w\in\mathbb{F}_{2}^{m}$, the size of
    $\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:D_{a}D_{b}F(x)=w\right\}$ (1)
    does not depend on $x\in\mathbb{F}_{2}^{n}$;
  2. 2.
    $F$ is plateaued with single amplitude if and only if the size of the set in eq. 1 does not depend on $x\in\mathbb{F}_{2}^{n}$, nor on $w\in\mathbb{F}_{2}^{m}$ when $w\neq 0$.
Theorem 4.3 is very useful for classifying plateaued functions, and in particular, we know that the size of the set
$\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:F(x)+F(x+a)+F(x+b)+F(x+a+b)=w\right\}$
is invariant of the choice of $x\in\mathbb{F}_{2}^{n}$ when $w\in\mathbb{F}_{2}^{n}$ is fixed and $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is a plateaued function. By replacing $b$ with $x+b$, the size of the above set is the same as
$|\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:F(x)+F(x+a)+F(b)+F(a+b)=w\right\}|$
$=\sum_{a\in\mathbb{F}_{2}^{n}}|\left\{b\in\mathbb{F}_{2}^{n}:F(x)+F(x+a)+F(b)+F(a+b)=w\right\}|$
$=\sum_{a\in\mathbb{F}_{2}^{n}}|\left\{b\in\mathbb{F}_{2}^{n}:D_{a}F(b)=F(x)+F(x+a)+w\right\}|$
$=\sum_{a\in\mathbb{F}_{2}^{n}}\delta_{F}(a,F(x)+F(x+a)+w).$
By replacing $a$ by $x+a$, we then have
$|\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:D_{a}D_{b}F(x)=w\right\}|=\sum_{a\in\mathbb{F}_{2}^{n}}\delta_{F}(a+x,F(a)+F(x)+w).$
In the particular case that $F$ is also APN, then $\delta_{F}=2\gamma_{F}+2^{n}\delta_{(0,0)}$. So when $F$ is APN and $w\neq 0$, we have $\sum_{a\in\mathbb{F}_{2}^{n}}\delta_{F}(a+x,F(a)+F(x)+w)=2\sum_{a\in\mathbb{F}_{2}^{n}}\gamma_{F}(a+x,F(a)+F(x)+w)$, and by Proposition 3.3 we know that $\sum_{a\in\mathbb{F}_{2}^{n}}\gamma_{F}(a+x,F(a)+F(x)+w)=3\operatorname{mult}_{\mathcal{G}_{F}}(x,F(x)+w)$. Hence
$|\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:D_{a}D_{b}F(x)=w\right\}|=6\operatorname{mult}_{\mathcal{G}_{F}}(x,F(x)+w)$ (2)
when $F$ is APN and $w\neq 0$. So Theorem 4.3 implies that if $F$ is plateaued APN and $w\neq 0$, then $\operatorname{mult}_{\mathcal{G}_{F}}(x,F(x)+w)$ does not depend on $x\in\mathbb{F}_{2}^{n}$. For $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ quadratic, it was shown in [5, Proposition 5] that $\Pi_{F}^{\beta}$ does not depend on $\beta\in\mathbb{F}_{2}^{n}$, and so if $F$ is APN quadratic, the exclude multiplicities $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ as $b$ ranges across $\mathbb{F}_{2}^{n}\setminus\left\{a\right\}$ does not depend on the choice of $a\in\mathbb{F}_{2}^{n}$ by Proposition 3.3. In the following proposition, we prove this result for all plateaued APN functions.
Proposition 4.4.
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be plateaued APN. If $a,b,c\in\mathbb{F}_{2}^{n}$ such that $b\neq F(a)$, then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\operatorname{mult}_{\mathcal{G}_{F}}(c,b+F(a)+F(c))$. In particular, the exclude multiplicities of $\mathcal{G}_{F}$ contained in $\left\{a\right\}\times(\mathbb{F}_{2}^{n}\setminus\left\{F(a)\right\})$ does not depend on the choice of $a$.
Proof.
For any $w\in\mathbb{F}_{2}^{n}\setminus\left\{0\right\}$, since $F$ is plateaued APN, we know from Theorem 4.3 and Equation 2 that the value of $\operatorname{mult}_{\mathcal{G}_{F}}(x,F(x)+w)$ does not depend on the choice of $x\in\mathbb{F}_{2}^{n}$. For all $a\in\mathbb{F}_{2}^{n}$, let
$X_{a}=\left\{a\right\}\times(\mathbb{F}_{2}^{n}\setminus\left\{F(a)\right\}),$
and define $\pi_{a,0}\colon X_{a}\to X_{0}$ to be the permutation given by $\pi_{a,0}(a,b)=(0,b+F(a)+F(0))$ for all $b\neq F(a)$. Then $\pi_{a,0}$ also preserves exclude multiplicities because for any $(a,b)\notin\mathcal{G}_{F}$ with $w=F(a)+b\neq 0$, we have
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ $=\operatorname{mult}_{\mathcal{G}_{F}}(a,F(a)+w)=\operatorname{mult}_{\mathcal{G}_{F}}(0,F(0)+w)=\operatorname{mult}_{\mathcal{G}_{F}}(0,F(0)+F(a)+b)$
$=\operatorname{mult}_{\mathcal{G}_{F}}(\pi_{a,0}(a,b)).$
Note that the inverse of $\pi_{a,0}$ is given by $(0,b)\mapsto(a,F(a)+F(0)+b)$. For all $a,c\in\mathbb{F}_{2}^{n}$, let $\pi_{a,c}\colon X_{a}\to X_{c}$ be the permutation $\pi_{a,c}=\pi_{c,0}^{-1}\circ\pi_{a,0}$. Then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\operatorname{mult}_{\mathcal{G}_{F}}(\pi_{a,c}(a,b))=\operatorname{mult}_{\mathcal{G}_{F}}(c,b+F(a)+F(c))$ for all $a,b,c\in\mathbb{F}_{2}^{n}$ with $b\neq F(a)$. ∎
We then immediately have the following divisibility condition on the frequencies of exclude multiplicities as a corollary of Proposition 4.4.
Corollary 4.5.
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued APN function. For any non-negative integer $k$, let $m_{k}=|\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=k\right\}|$. Then $2^{n}|m_{k}$ for all $k\geq 0$.
Proof.
For any non-negative integer $k$, let $z_{k}=|\left\{b\in\mathbb{F}_{2}^{n}\setminus\left\{F(0)\right\}:\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=k\right\}|$. By Proposition 4.4, we know that for any $a\in\mathbb{F}_{2}^{n}$, there exists a bijection from $\left\{a\right\}\times(\mathbb{F}_{2}^{n}\setminus\left\{F(a)\right\})$ to $\left\{0\right\}\times(\mathbb{F}_{2}^{n}\setminus\left\{F(0)\right\})$ that preserves exclude multiplicity. Therefore, $m_{k}=2^{n}z_{k}$ for all $k\geq 0$, so $2^{n}|m_{k}$. ∎
From this, we prove that the Brinkmann-Leander-Edel-Pott function $F\colon\mathbb{F}_{2^{6}}\to\mathbb{F}_{2^{6}}$ (see [23], [11, Section 11.5.3], or [4]), given by
$F(x)$ $=x^{3}+u^{17}(x^{17}+x^{18}+x^{20}+x^{24})+u^{14}(u^{18}x^{9}+u^{36}x^{18}+u^{9}x^{36}+x^{21}+x^{42})$
$+u^{14}\operatorname{Tr}(u^{52}x^{3}+u^{6}x^{5}+u^{19}x^{7}+u^{28}x^{11}+u^{2}x^{13}),$
where $u\in\mathbb{F}_{2^{n}}$ is primitive, cannot be CCZ-equivalent to a plateaued function. This improves our understanding on this sporadic APN function as the Brinkmann-Leander-Edel-Pott function was only previously known to not be CCZ-equivalent to a quadratic function or monomial function [23].
Corollary 4.6.
The Brinkmann-Leander-Edel-Pott function is not CCZ-equivalent to any plateaued APN function.
Proof.
Let $F\colon\mathbb{F}_{2^{6}}\to\mathbb{F}_{2^{6}}$ be the Brinkmann-Leander-Edel-Pott function. We have computed that $\mathcal{G}_{F}$ has points of exclude multiplicity $5,7,9,11,13$, and $15$ with frequencies $40,360,1296,1616,648$, and $72$, respectively. By Corollary 4.5 and the fact that exclude multiplicities are preserved under CCZ-equivalence, any function CCZ-equivalent to a plateaued function must satisfy the property that the frequency of any exclude multiplicity must be divisible by $2^{n}$. Hence, the Brinkmann-Leander-Edel-Pott function is not CCZ-equivalent to any plateaued APN function as $2^{6}$ does not divide $40$ (or any other frequency listed above). ∎
Our proof technique for showing that the Brinkmann-Leander-Edel-Pott function is not CCZ-equivalent to a plateaued function also yields a practical test that can be applied to newly discovered APN functions in the future.
We can also describe the exclude multiplicities of $\mathcal{G}_{F}$ more directly in terms of the amplitudes of the component functions of $F$.
Proposition 4.7.
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued APN function, and let $\lambda_{v}$ denote the amplitude of $v\cdot F$. Then for any $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, we have
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\frac{1}{6\cdot 2^{n}}\sum_{v\in\mathbb{F}_{2}^{n}}(-1)^{v\cdot(F(a)+b)}\lambda_{v}^{2}.$ (3)
Proof.
For any $(u,v)\in(\mathbb{F}_{2}^{n})^{2}\notin\mathcal{G}_{F}$, we have $W_{F}^{3}(u,v)=\lambda_{v}^{2}W_{F}(u,v)$ as $F$ is plateaued. Recall that for any $(a,b)\notin\mathcal{G}_{F}$, we have $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\frac{1}{6\cdot 2^{2n}}\widehat{W_{F}^{3}}(a,b)$, and by Proposition 4.4, we also have that $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\operatorname{mult}_{\mathcal{G}_{F}}(0,b+F(a)+F(0))$. So, for $(a,b)\notin\mathcal{G}_{F}$,
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ $=\frac{1}{6\cdot 2^{2n}}\sum_{(u,v)\in(\mathbb{F}_{2}^{n})^{2}}(-1)^{v\cdot(b+F(a)+F(0))}\lambda_{v}^{2}W_{F}(u,v)$
$=\frac{1}{6\cdot 2^{2n}}\sum_{v\in\mathbb{F}_{2}^{n}}(-1)^{v\cdot(b+F(a)+F(0))}\lambda_{v}^{2}\sum_{x\in\mathbb{F}_{2}^{n}}(-1)^{v\cdot F(x)}\sum_{u\in\mathbb{F}_{2}^{n}}(-1)^{x\cdot u}$
$=\frac{1}{6\cdot 2^{n}}\sum_{v\in\mathbb{F}_{2}^{n}}(-1)^{v\cdot(b+F(a))}\lambda_{v}^{2}.$
As seen in Proposition 4.7, the amplitudes of the component functions of an APN plateaued function can be used to describe the exclude multiplicities of its graph. For $c\neq 0$, it is clear that we have the equalities
$2\sum_{v\in\left\{c\right\}^{\perp}}\lambda_{v}^{2}-\sum_{v\in\mathbb{F}_{2}^{n}}\lambda_{v}^{2}=\sum_{v\in\mathbb{F}_{2}^{n}}(-1)^{v\cdot c}\lambda_{v}^{2}=\sum_{v\in\mathbb{F}_{2}^{n}}\lambda_{v}^{2}-2\sum_{v\notin\left\{c\right\}^{\perp}}\lambda_{v}^{2}.$ (4)
By using the following proposition, we will be able to rephrase minimizing exclude multiplicity in terms of minimizing the sum of squared amplitudes in any linear hyperplane as the sum $\sum_{v\in\mathbb{F}_{2}^{n}}\lambda_{v}^{2}$ is equal to $2^{n}(3\cdot 2^{n}-2)$ precisely when $F$ is APN.
Proposition 4.8 ([10, Proposition 9]).
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be an APN plateaued function, and let $\lambda_{v}$ denote the amplitude of $v\cdot F$ for all $v\in\mathbb{F}_{2}^{n}$. Then $F$ is APN if and only if
$\sum_{v\in\mathbb{F}_{2}^{n}}\lambda_{v}^{2}\leq 2^{n}(3\cdot 2^{n}-2).$ (5)
In particular, if $F$ is APN, then eq. 5 is an equality.
Corollary 4.9.
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be an APN plateaued function, and let $\lambda_{v}$ denote the amplitude of $v\cdot F$ for all $v\in\mathbb{F}_{2}^{n}$. Then for any $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, we have
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\frac{1}{3\cdot 2^{n}}\left(\sum_{v\in\left\{b+F(a)\right\}^{\perp}}\lambda_{v}^{2}-2^{n-1}(3\cdot 2^{n}-2)\right).$
Proof.
Let $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$. By eq. 3, we have
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\frac{1}{6\cdot 2^{n}}\sum_{v\in\mathbb{F}_{2}^{n}}(-1)^{v\cdot(F(a)+b)}\lambda_{v}^{2}=\frac{1}{6\cdot 2^{n}}\left(2\sum_{v\in\left\{b+F(a)\right\}^{\perp}}\lambda_{v}^{2}-\sum_{v\in\mathbb{F}_{2}^{n}}\lambda_{v}^{2}\right).$
The result then follows by applying Proposition 4.8. ∎
Corollary 4.10.
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued APN function, and let $\lambda_{v}$ denote the amplitude of $v\cdot F$ for all $v\in\mathbb{F}_{2}^{n}$. If for a non-negative integer $k$, we have
$\sum_{v\in\left\{c\right\}^{\perp}}\lambda_{v}^{2}\geq 3\cdot 2^{n}k+2^{n-1}(3\cdot 2^{n}-2)$ (6)
for all nonzero $c\in\mathbb{F}_{2}^{n}$, then $e_{\min}(\mathcal{G}_{F})\geq k$.
Proof.
Suppose $k$ is a non-negative integer such that for all nonzero $c\in\mathbb{F}_{2}^{n}$ eq. 6 holds. For any $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, we have by Corollary 4.9 that
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ $\geq\frac{1}{3\cdot 2^{n}}\left(3\cdot 2^{n}k+2^{n-1}(3\cdot 2^{n}-2)-2^{n-1}(3\cdot 2^{n}-2)\right)=k.$
So $e_{\min}(\mathcal{G}_{F})\geq k$. ∎
Although Corollary 4.10 follows easily from previous statements, it helps shape our approach on finding a lower bound on the minimum exclude multiplicity of the graph of a plateaued APN function. By using a geometric constraint on the structure of ${\mathcal{B}({F})}=\left\{v\in\mathbb{F}_{2}^{n}:v\cdot F\text{ is bent}\right\}$, we can determine a lower bound between plateaued APN functions and other APN functions. We use the following corollary which follows from a classical result of [3].
Corollary 4.11 ([34, Corollary 1]).
A set $S\subseteq(\mathbb{F}_{2}^{n}\setminus\left\{0\right\})$ that intersects every (n+1k)(n+1-k)-dimensional subspace of $\mathbb{F}_{2}^{n}$ has at least $2^{k}-1$ elements with equality if and only if $S\cup\left\{0\right\}$ is a $k$-dimensional subspace of $\mathbb{F}_{2}^{n}$.
The following lemma is a simple generalization of Theorem 3 from [34].
Lemma 4.12.
Assume $n$ is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be any vectorial Boolean function. Let $S\subseteq\mathbb{F}_{2}^{n}$ be a linear subspace of dimension $d$ where $\frac{n}{2}+1\leq d\leq n$. Then $|S\setminus{\mathcal{B}({F})}|\geq 2^{d-\frac{n}{2}}$ with equality if and only if $S\setminus{\mathcal{B}({F})}$ is a $(d-\frac{n}{2})$-dimensional linear subspace of $S$. Hence, there are at least $2^{d-\frac{n}{2}}-1$ nonzero, non-bent component functions $v\cdot F$ with $v\in S$.
Proof.
In [34, Proof of Theorem 3], it was shown that there cannot exist a $(\frac{n}{2}+1)$-dimensional linear subspace $T$ of $\mathbb{F}_{2}^{n}$ such that $T\setminus\left\{0\right\}$ is contained in $\mathcal{B}({F})$} (otherwise we could easily construct a function $\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n/2+1}$ whose component functions are all bent, but this is impossible by Nyberg’s bound). Hence, $S\cap{\mathcal{B}({F})}$ cannot contain all nonzero elements of a linear subspace of dimension $\frac{n}{2}+1$ because it is a subset of $\mathcal{B}({F})$}. So $S\setminus({\mathcal{B}({F})}\cup\left\{0\right\})$ must intersect every linear subspace of $S$ of dimension $\frac{n}{2}+1$. Also, note that $\frac{n}{2}+1=d+1-(d-\frac{n}{2})$. So, by applying a linear isomorphism from $S$ to $\mathbb{F}_{2}^{d}$ and considering the image of $S\setminus({\mathcal{B}({F})}\cup\left\{0\right\})$, it follows from Corollary 4.11 that
$|S\setminus({\mathcal{B}({F})}\cup\left\{0\right\})|\geq 2^{d-\frac{n}{2}}-1$
with equality if and only if $S\setminus{\mathcal{B}({F})}$ is a $(d-\frac{n}{2})$-dimensional subspace of $S$. ∎
From this, we are able to derive lower bounds on the exclude multiplicities of the graphs of APN functions (for $n$ even).
Theorem 4.13.
Assume $n\geq 4$ is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued APN function. Then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq 2^{\frac{n}{2}-1}-1$ for all $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$.
Proof.
Let $c=F(a)+b$ for $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$. Let $N=|\left\{c\right\}^{\perp}\setminus({\mathcal{B}({F})}\cup\left\{0\right\})|$. By applying Corollary 4.9, we have
$\sum_{v\in\left\{c\right\}^{\perp}}\lambda_{v}^{2}$ $=2^{2n}+(2^{n-1}-1-N)2^{n}+\sum_{v\in\left\{c\right\}^{\perp}\setminus({\mathcal{B}({F})}\cup\left\{0\right\})}\lambda_{v}^{2}$
$\geq 2^{2n}+(2^{n-1}-1-N)2^{n}+N2^{n+2}$
$=3\cdot 2^{n}N+2^{n-1}(3\cdot 2^{n}-2).$
By Lemma 4.12, we have $N\geq 2^{\frac{n}{2}-1}-1$, so $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq e_{\min}(\mathcal{G}_{F})\geq N\geq 2^{\frac{n}{2}-1}-1$ by Corollary 4.10. ∎
Our lower bound in Theorem 4.13 is a dramatic increase over the previously known $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq 1$ when $F$ is plateaued APN. It would be interesting to further improve this lower bound or find an example of a family of plateaued APN functions for which it is tight. The lower bound on the distance from plateaued APN functions to other APN functions immediately follows.
Proof of Theorem 1.1.
Apply Lemma 3.1 to Theorem 4.13. ∎
When $n$ is odd, any plateaued APN function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is AB, and so we know by a result of [35] that $\mathcal{G}_{F}$ is a maximal Sidon set with all points outside of the graph having exclude multiplicity $\frac{2^{n}-2}{6}$ which is odd (see Section 4.1). However, when $n$ is even, the exclude multiplicities of $\mathcal{G}_{F}$ are much more unclear. Computational results seem to suggest that if $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is a plateaued APN function, then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd for all $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, and we conjecture in 4.15 that this is true. In the following proposition, we describe the parity of the exclude multiplicities of $\mathcal{G}_{F}$ for $F$ plateaued APN in terms of the Fourier transform of the indicator function $1_{\mathcal{B}({F})}\colon\mathbb{F}_{2}^{n}\to\left\{0,1\right\}$ of $\mathcal{B}({F})$}.
Proposition 4.14.
Assume $n$ is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be plateaued APN function and let $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$. Then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd if and only if $\widehat{1_{\mathcal{B}({F})}}(F(a)+b)\equiv 2\mod 4$. Equivalently, $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd if and only if $|{\mathcal{B}({F})}\cap\left\{F(a)+b\right\}^{\perp}|$ is even.
Proof.
For all $v\in\mathbb{F}_{2}^{n}$, let $\lambda_{v}$ denote the amplitude of $v\cdot F$, and let $m_{v}$ be the non-negative integer such that $\lambda_{v}^{2}=2^{n+m_{v}}$. Note that $m_{v}$ is even and $0\leq m_{v}\leq n$ for all $v\in\mathbb{F}_{2}^{n}$. Let $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, and let $c=F(a)+b$. By Proposition 4.7, we have
$6\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\frac{1}{2^{n}}\sum_{v\in\mathbb{F}_{2}^{n}}(-1)^{v\cdot c}2^{n}2^{m_{v}}=\sum_{v\in\mathbb{F}_{2}^{n}}(-1)^{v\cdot c}2^{m_{v}}.$
Since $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is an integer, we then know that $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd if and only if
$\sum_{v\in\mathbb{F}_{2}^{n}}(-1)^{v\cdot c}2^{m_{v}}\equiv 6\mod 12.$
Also, $2^{m_{v}}\equiv 1\mod 12$ if $m_{v}=0$ (i.e. $v\cdot F$ is bent) and $2^{m_{v}}\equiv 4\mod 12$ if $m_{v}\geq 2$ as $m_{v}$ is even. For $v\in\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}$, we then can write $2^{m_{v}}=12r_{v}+4$ for some non-negative integer $r_{v}$, so $r_{v}=\frac{2^{m_{v}}-4}{12}$. Hence
$6\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ $=\sum_{v\in{\mathcal{B}({F})}}(-1)^{v\cdot c}+\sum_{v\in\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}}(-1)^{v\cdot c}(12r_{v}+4)$
$=\widehat{1_{\mathcal{B}({F})}}(c)+4\cdot\widehat{1_{\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}}}(c)+12\sum_{v\in\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}}(-1)^{v\cdot c}r_{v}.$
So $6\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\equiv\widehat{1_{\mathcal{B}({F})}}(c)+4\cdot\widehat{1_{\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}}}(c)\mod 12$. Note that $\widehat{1_{\mathcal{B}({F})}}(c)+4\cdot\widehat{1_{\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}}}(c)\equiv 6\mod 12$ if and only if $\widehat{1_{\mathcal{B}({F})}}(c)\equiv 2\mod 4$ and $\widehat{1_{\mathcal{B}({F})}}(c)+\widehat{1_{\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}}}(c)\equiv 0\mod 3$. However, we already know that $\widehat{1_{\mathcal{B}({F})}}(c)+\widehat{1_{\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}}}(c)=0$ as $c\neq 0$. So $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd if and only if $\widehat{1_{\mathcal{B}({F})}}(c)\equiv 2\mod 4$, and this proves the first part of the statement.
For $i\in\left\{0,1\right\}$, let $B_{i}=|\left\{v\in{\mathcal{B}({F})}:v\cdot c=i\right\}|$. Then $\widehat{1_{\mathcal{B}({F})}}(c)=|B_{0}|-|B_{1}|=|{\mathcal{B}({F})}|-2|B_{1}|$. It was shown in [30, Proposition 6.5] that the number of bent components of any plateaued function over an even dimension is $2\mod 4$, and so $|{\mathcal{B}({F})}|\equiv 2\mod 4$. Therefore, $\widehat{1_{\mathcal{B}({F})}}(c)\equiv 2-2|B_{1}|\mod 4$. So $\widehat{1_{\mathcal{B}({F})}}(c)\equiv 2\mod 4$ if and only if $|B_{1}|$ is even. Since the parity of $|B_{0}|$ and $|B_{1}|$ is the same, the second part of the result follows. ∎
Conjecture 4.15.
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be an APN plateaued function. Then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd for all $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$.
In Section 4.4, we prove the above conjecture in the affirmative in the case of $F$ being quadratic APN, but we were unable to find a proof for all plateaued APN functions.
We now express the exclude multiplicities of the graph of an APN function in terms of $\widehat{W_{F}^{4}}$.
Proposition 4.16.
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued APN function. For any $(a,b)\notin\mathcal{G}_{F}$, we have
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\frac{1}{6\cdot 2^{3n}}\widehat{W_{F}^{4}}(0,b+F(a)+F(0)).$
As a consequence, if $F(0)=0$ and $c\neq 0$, then $2^{n}\widehat{W_{F}^{3}}(0,c)=\widehat{W_{F}^{4}}(0,c)$.
Proof.
Let $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, and let $c=b+F(a)$. Notice that
$\frac{1}{2^{2n}}\widehat{W_{F}^{4}}(0,c)$ $=|\left\{(x,y,z,w)\in(\mathbb{F}_{2}^{n})^{4}:(x+y+z+w,F(x)+F(y)+F(z)+F(w))=(0,c)\right\}|$
$=|\left\{(x,y,z)\in(\mathbb{F}_{2}^{n})^{3}:F(x)+F(y)+F(z)+F(x+y+z)=(0,c)\right\}|$
$=|\left\{(x,a,b)\in(\mathbb{F}_{2}^{n})^{3}:F(x)+F(x+a)+F(x+b)+F(x+a+b)=c\right\}|$
$=\sum_{x\in\mathbb{F}_{2}^{n}}|\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:D_{a}D_{b}F(x)=c\right\}|.$
By Theorem 4.3 and eq. 2, we then know that $\frac{1}{6\cdot 2^{3n}}\widehat{W_{F}^{4}}(0,c)=\operatorname{mult}_{\mathcal{G}_{F}}(0,F(0)+c)$. The equation then follows as $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\operatorname{mult}_{\mathcal{G}_{F}}(0,b+F(a)+F(0))$ by Proposition 4.4. In the case that $F(0)=0$, then $2^{n}\widehat{W_{F}^{3}}(0,c)=\widehat{W_{F}^{4}}(0,c)$ for all nonzero $c\in\mathbb{F}_{2}^{n}$ as $\operatorname{mult}_{\mathcal{G}_{F}}(0,c)=\frac{1}{6\cdot 2^{2n}}\widehat{W_{F}^{3}}(0,c)$. ∎
From the above result, we can also express the exclude multiplicities of $\mathcal{G}_{F}$ in terms of the derivatives of $\gamma_{F}$. In Section 7, we will see why this is relevant and how it is connected to the existence of nontrivial linear structures of $\gamma_{F}$. In [13, Section 6], it was shown that for any APN function $F$, we have $W_{F}^{2}=2^{2n}\delta_{(0,0)}-W_{\gamma_{F}}+2^{n}$, (where $\delta_{(0,0)}$ takes value $1$ at $(0,0)$ and 0 elsewhere). So
$W_{F}^{4}=\delta_{(0,0)}(2^{4n}-2^{2n+1}W_{\gamma_{F}}+2^{3n+1})+W_{\gamma_{F}}^{2}-2^{n+1}W_{\gamma_{F}}+2^{2n}.$
Since $F$ is APN, the Hamming weight $\operatorname{wt}(\gamma_{F})=\sum_{(a,b)\in(\mathbb{F}_{2}^{n})^{2}}\gamma_{F}(a,b)$ of $\gamma_{F}$ is equal to $\binom{2^{n}}{2}$, see [14]. Therefore $W_{\gamma_{F}}(0,0)=2^{2n}-2\operatorname{wt}(\gamma_{F})=2^{n}$. Applying the Fourier-Hadamard transform to both sides of the equation above, we have
$\widehat{W_{F}^{4}}$ $=2^{4n}-2^{2n+1}W_{\gamma_{F}}(0,0)+2^{3n+1}+\widehat{W_{\gamma_{F}}^{2}}-2^{n+1}\widehat{W_{\gamma_{F}}}+2^{4n}\delta_{(0,0)}$
$=2^{4n}-2^{3n+1}+2^{3n+1}+\widehat{W_{\gamma_{F}}^{2}}-2^{n+1}\widehat{W_{\gamma_{F}}}+2^{4n}\delta_{(0,0)}$
$=2^{4n}+2^{4n}\delta_{(0,0)}+\widehat{W_{\gamma_{F}}^{2}}-2^{n+1}\widehat{W_{\gamma_{F}}}$.
Observe that $\widehat{W_{\gamma_{F}}}=2^{2n}+2^{3n}\delta_{(0,0)}-\widehat{W_{F}^{2}}=2^{2n}+2^{3n}\delta_{(0,0)}-2^{2n}\delta_{F}$. So $\widehat{W_{\gamma_{F}}}(0,b)=2^{2n}$ for all $b\neq 0$. Therefore, for all $b\neq 0$, we have $\widehat{W_{F}^{4}}(0,b)=2^{4n}+\widehat{W_{\gamma_{F}}^{2}}-2^{3n+1}$, so
$\widehat{W_{F}^{4}}(0,b)=2^{3n}(2^{n}-2)+2^{2n}\Delta_{\gamma_{F}}(0,b)$ (7)
as the autocorrelation $\Delta_{\gamma_{F}}(u,v)=\sum_{(x,y)\in(\mathbb{F}_{2}^{n})^{2}}(-1)^{\gamma_{F}(x,y)+\gamma_{F}(x+u,y+v)}=2^{2n}-2\operatorname{wt}(D_{(u,v)}\gamma_{F})$ of $\gamma_{F}$ is equal to $\frac{1}{2^{2n}}\widehat{W_{\gamma_{F}}^{2}}$ (cf. [11]). We then have the following proposition.
Proposition 4.17.
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued APN function. Then for all $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, we have the equality
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\frac{2^{n}-2}{6}+\frac{\Delta_{\gamma_{F}}(0,b+F(a))}{6\cdot 2^{n}}.$
Proof.
Let $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$. By Proposition 4.4, we have $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\operatorname{mult}_{\mathcal{G}_{F}}(0,b+F(a))$. Therefore, by applying Proposition 4.16 and eq. 7, we have
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ $=\operatorname{mult}_{\mathcal{G}_{F}}(0,b+F(0)+F(a))$
$=\frac{1}{6\cdot 2^{3n}}\widehat{W_{F}^{4}}(0,b+F(a))$
$=\frac{2^{n}-2}{6}+\frac{\Delta_{\gamma_{F}}(0,b+F(a))}{6\cdot 2^{n}}.$
4.3 3-to-1 plateaued functions
Throughout this subsection, we assume that $n$ is even. We say a function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is 33-to-11 if every point in the image of $F$ has a preimage of size $3$ except a single point. Note that all 3-to-1 plateaued functions are necessarily APN (see [29, Theorem 12]).
A subset $D\subseteq G$ of an additive group $G$ is a partial difference set with parameters $(v,k,\lambda,\mu)$ if every non-identity element of $D$ (resp. $G\setminus D$) can be written as xyx-y with distinct $x,y\in S$ in exactly $\lambda$ (resp. $\mu$) ways. The following theorem was originally only proved for crooked 33-to-11 functions (in short, those functions such that $\operatorname{Im}D_{a}F$ is an affine hyperplane for all $a\neq 0$), but its proof only relied on the fact that all crooked functions are plateaued, so we write it for all 33-to-11 plateaued functions.
Theorem 4.18 ([7, Theorem 4]).
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued such that $F(0)=0$ and $F$ is 3-to-1 on $\mathbb{F}_{2}^{n}\setminus\left\{0\right\}$. Then $\operatorname{Im}F\setminus\left\{0\right\}$ is a partial difference set, with parameters $(2^{n},\frac{2^{n}-1}{3},\frac{2}{3}(\alpha(n)-1),\frac{2}{3}\beta(n))$, where $\alpha(n)=\frac{2^{n}+(-2)^{\frac{n}{2}+1}-2}{6}$ and $\beta(n)=\frac{2^{n}+(-2)^{\frac{n}{2}}-2}{6}$.
Moreover, [7, Corollary 4] describes the multiplicities of the nonzero elements of the multiset $\left\{F(x)+F(y)+F(x+y):x,y\in\mathbb{F}_{2}^{n}\right\}$ when $F$ is quadratic 33-to-11, and indeed, this can be easily generalized to all plateaued 3-to-1 APN functions.
Corollary 4.19.
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued 33-to-11 function with $F(0)=0$ such that $\operatorname{Im}F\setminus\left\{0\right\}$ is a partial difference set with parameters $(2^{n},\frac{2^{n}-1}{3},\frac{2}{3}(\alpha(n)-1),\frac{2}{3}\beta(n))$. Let $b\in\mathbb{F}_{2}^{n}\setminus\left\{0\right\}$. Then $\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=\alpha(n)$ if $b\in\operatorname{Im}F$ and $\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=\beta(n)$ if $b\notin\operatorname{Im}F$.
These two results above (in terms of quadratic 3-to-1 functions) and [5, Proposition 5] (which states that $\Pi_{F}^{\beta}$ does not depend on $\beta\in\mathbb{F}_{2}^{n}$ when $F$ is quadratic) were used in [7] to determine the exact exclude multiplicities of $\mathcal{G}_{F}$ and their frequencies in the case of $F$ being a quadratic 33-to-11 function. This can be done because any plateaued function $F$ with all component functions unbalanced satisfies
$|\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:D_{a}D_{b}F(x)=w\right\}|=|\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:F(a)+F(b)=w\right\}|$ (8)
for all $x,w\in\mathbb{F}_{2}^{n}$ (see [10, Theorem 2]). Therefore, if $F$ is plateaued APN with all unbalanced components and $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, then
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\frac{1}{6}|\left\{(u,v)\in(\mathbb{F}_{2}^{n})^{2}:F(u)+F(v)=F(a)+b\right\}|$
by combining eqs. 2 and 8. In Proposition 4.4, we proved that the exclude multiplicities of the graph of a plateaued APN function $F$ are completely determined by the exclude multiplicities of points in $\left\{0\right\}\times(\mathbb{F}_{2}^{n}\setminus\left\{0\right\})$. We use this to prove the following result.
Theorem 4.20.
Suppose $n\geq 4$ is even, and let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued 33-to-11 function. Then
  1. 1.
    there are $2^{n}\cdot\frac{2^{n}-1}{3}$ exclude points of $\mathcal{G}_{F}$ with multiplicity $\alpha(n)$;
  2. 2.
    there are $2^{n+1}\cdot\frac{2^{n}-1}{3}$ exclude points of $\mathcal{G}_{F}$ with multiplicity $\beta(n)$;
  3. 3.
    every point in $(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$ has exclude multiplicity $\alpha(n)$ or $\beta(n)$.
Proof.
Without loss of generality, assume $F(0)=0$. By Theorem 4.18, $\operatorname{Im}F\setminus\left\{0\right\}$ is a partial difference set with parameters $(2^{n},\frac{2^{n}-1}{3},\frac{2}{3}(\alpha(n)-1),\frac{2}{3}\beta(n))$. The result then immediately follows by applying Proposition 4.4 and Corollary 4.19 (cf. [7]) and using the same proof as [7, Corollary 4]. ∎
We now have generalized the lower bounds on the distances of APN functions to quadratic 3-to-1 functions to plateaued 3-to-1 functions, which also includes Kasami functions as shown by Yoshiara [36]. Theorem 1.2 immediately follows.
Proof of Theorem 1.2.
Apply Lemma 3.1 to Theorem 4.20. ∎
4.4 Quadratic APN functions
Recall that $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is quadratic if every derivative $D_{a}F(x)=F(x)+F(x+a)$ of $F$ is an affine function. Quadratic APN functions are particularly well-behaved as the images of their derivatives are affine hyperplanes (more generally, this property is called crookedness, see [11, 31]). Moreover, quadratic APN functions are equipped with a unique so-called ortho-derivative.
Definition 4.21.
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a quadratic APN function. The ortho-derivative of $F$ is the unique function $\pi_{F}\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ such that $\pi_{F}(0)=0$ and $\pi_{F}(a)\neq 0$ for all $a\in\mathbb{F}_{2}^{n}\setminus\left\{0\right\}$ and
$\pi_{F}(a)\cdot(F(x)+F(x+a)+F(0)+F(a))=0$
for all $a,x\in\mathbb{F}_{2}^{n}$.
The ortho-derivative of an APN quadratic function $F$ is the function $\pi_{F}$ such that $\pi_{F}(a)=0$ and $\left\{\pi_{F}(a)\right\}^{\perp}$ is the linear part of the affine hyperplane $\operatorname{Im}D_{a}F$ for all nonzero $a\in\mathbb{F}_{2}^{n}$. In [18], properties of the ortho-derivative were studied, and in particular, there exists a connection between the ortho-derivative of $F$ and exclude multiplicity.
Proposition 4.22 ([18]).
Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a quadratic APN function, and let $\pi_{F}\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be its nontrivial ortho-derivative. Then for any $b\in\mathbb{F}_{2}^{n}\setminus\left\{0\right\}$,
$\frac{1}{2^{2n}}\widehat{W_{F}^{4}}(0,b)=2^{n+1}(2^{n}-1-\operatorname{wt}(b\cdot\pi_{F})).$
Therefore if $F$ is a quadratic APN function with $F(0)=0$ and $b\neq 0$, then
$\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=\frac{1}{6\cdot 2^{2n}}\widehat{W_{F}^{3}}(0,b)=\frac{1}{6\cdot 2^{3n}}\widehat{W_{F}^{4}}(0,b)=\frac{1}{3}(2^{n}-1-\operatorname{wt}(b\cdot\pi_{F})).$ (9)
Remark 4.23.
Let $F$ be a quadratic APN function with $F(0)=0$. If $n$ is odd, then all nonzero components $b\cdot\pi_{F}$ of the ortho-derivative of $F$ are balanced, i.e. $\pi_{F}$ is a permutation (this was observed in [18]). This is because when $n$ is odd, $F$ is AB implying $\operatorname{wt}(b\cdot\pi_{F})=2^{n}-1-3\cdot\frac{2^{n-1}-1}{3}=2^{n-1}$ by Theorem 4.1. If $n$ is even, then $\operatorname{wt}(b\cdot\pi_{F})$ is a multiple of $3$ for all $b\neq 0$ as $3$ divides $2^{n}-1-\operatorname{wt}(b\cdot\pi_{F})$ and $2^{n}\equiv 1\mod 3$.
Note that eq. 9 tells us that for $F$ APN quadratic, obtaining a lower bound on the exclude points of $\mathcal{G}_{F}$ is equivalent to finding an upper bound on the weights of the components of $\pi_{F}$. By applying Theorem 4.13, we then obtain an upper bound on the weights of the component functions of APN quadratic functions.
Proposition 4.24.
Assume $n\geq 4$ is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a quadratic APN function with $F(0)=0$. For any $b\neq 0$, the weight of the component function $b\cdot\pi_{F}$ of the ortho-derivative of $F$ satisfies $\operatorname{wt}(b\cdot\pi_{F})\leq 2^{n}+2-3\cdot 2^{\frac{n}{2}-1}$.
Proof.
By Theorem 4.13, we have that $\operatorname{mult}_{\mathcal{G}_{F}}(0,b)\geq 2^{\frac{n}{2}-1}-1$ for all $b\neq 0$. The result then follows by applying this bound to eq. 9. ∎
It was also shown in [18] that the ortho-derivative of a quadratic APN function has algebraic degree at most n2n-2. For any function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$, it is well-known that $F$ has algebraic degree $n$ if and only if $\sum_{x\in\mathbb{F}_{2}^{n}}F(x)\neq 0$. Therefore, if $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is quadratic APN, then $\sum_{x\in\mathbb{F}_{2}^{n}}\pi_{F}(x)=0$, implying that all component functions of $\pi_{F}$ have even Hamming weight as for any $v\neq 0$, we have $\operatorname{wt}(v\cdot\pi_{F})\equiv\sum_{x\in\mathbb{F}_{2}^{n}}v\cdot\pi_{F}(x)\pmod{2}=v\cdot\sum_{x\in\mathbb{F}_{2}^{n}}\pi_{F}(x)=0.$ This allows us to prove that the exclude multiplicities of the graph of an APN quadratic function are always odd.
Proposition 4.25.
Assume $n\geq 3$. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a quadratic APN function. Then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd for all $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$. Hence, $|\mathcal{B}(F)\cap H|$ is even for any affine hyperplane $H$ of $\mathbb{F}_{2}^{n}$.
Proof.
If $n$ is odd, then $F$ is AB and $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\frac{2^{n}-2}{6}$ for all $(a,b)\notin\mathcal{G}_{F}$. So, assume $n$ is even, and without loss of generality, assume $F(0)=0$. As mentioned above, the algebraic degree of $\pi_{F}$ is at most n2n-2, implying every component function of $\pi_{F}$ has even Hamming weight. Therefore, by eq. 9, for any $b\neq 0$, we know that $\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=\frac{1}{3}(2^{n}-1-\operatorname{wt}(b\cdot\pi_{F}))$ is odd. Thus, by applying Proposition 4.4, we know that any exclude point of $\mathcal{G}_{F}$ has odd multiplicity. The parity of $|\mathcal{B}(F)\cap H|$ must then be even for any affine hyperplane $H$ of $\mathbb{F}_{2}^{n}$, see Proposition 4.14. ∎
Note that it does not always hold that the exclude multiplicities of the graph of an APN function must be odd. For example, when $n=5$, the graph of the APN inverse function has exclude points of multiplicity $3,4,5$, and $6$.
As a corollary of Proposition 4.25, every quadratic APN function has a maximal Sidon set as its graph, and as previously mentioned, this is already known since all quadratic functions are plateaued. Nonetheless, it is interesting that Proposition 4.25 now provides a fourth method (the other three being the work of [6], the usage of the D-property in [12], and Theorem 4.13) for proving no quadratic APN function has a non-maximal Sidon set as its graph.
5 Power functions
In this subsection, we consider the case when $F\colon\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ is of the form $F(x)=x^{d}$ for some integer $d$. If $F$ is an APN power function, then we will see that the possible values of $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ does not depend on $a\in\mathbb{F}_{2^{n}}^{\ast}$ as $b$ ranges across $\mathbb{F}_{2^{n}}$ (i.e. $\Pi_{F}^{\beta}$ does not depend on $\beta\in\mathbb{F}_{2^{n}}^{\ast}$). Also, we will fully determine $\operatorname{mult}_{\mathcal{G}_{F}}(0,a)$ and $\operatorname{mult}_{\mathcal{G}_{F}}(a,0)$ when $F$ is APN, $n$ is odd, and $a\neq 0$.
First, recall that for any power function $F(x)=x^{d}$, we have $W_{F}(u,v)=W_{F}(1,u^{-d}v)$ for all $u\neq 0$, and in the case when $F$ is a permutation, we then have for $v\neq 0$ that $W_{F}(u,v)=W_{F}(uv^{-\frac{1}{d}},1)$ where $\frac{1}{d}$ is the inverse of $d$ modulo $2^{n}-1$ [11]. The first of these two properties allows us to prove the following result.
Proposition 5.1.
Let $F\colon\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ be an APN power function $F(x)=x^{d}$. If $a,b,c\in\mathbb{F}_{2^{n}}$ such that $a,c\neq 0$ and $b\neq a^{d}$, then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\operatorname{mult}_{\mathcal{G}_{F}}(c,(a^{-1}c)^{d}b)$. In particular, the exclude multiplicities of $\mathcal{G}_{F}$ contained in $\left\{a\right\}\times(\mathbb{F}_{2^{n}}\setminus\left\{a\right\})$ does not depend on the choice of $a\in\mathbb{F}_{2^{n}}^{\ast}$.
Proof.
Let $a,b\in\mathbb{F}_{2^{n}}$ such that $a\neq 0$ and $b\neq a^{d}$. Recall that we have the equality $W_{F}(u,v)=W_{F}(1,u^{-d}v)$ for all $u,v\in\mathbb{F}_{2}^{n}$ where $u\neq 0$. So
$\widehat{W_{F}^{3}}(a,b)$ $=\sum_{(u,v)\in\mathbb{F}_{2^{n}}^{2}}(-1)^{\operatorname{Tr}(ua)+\operatorname{Tr}(vb)}W_{F}^{3}(u,v)$
$=\sum_{\begin{subarray}{c}(u,v)\in\mathbb{F}_{2^{n}}^{2}\\ u\neq 0\end{subarray}}(-1)^{\operatorname{Tr}(u)+\operatorname{Tr}(vb)}W_{F}^{3}(ua^{-1},v)+\sum_{v\in\mathbb{F}_{2^{n}}}(-1)^{\operatorname{Tr}(vb)}W_{F}^{3}(0,v)$
$=\sum_{\begin{subarray}{c}(u,v)\in\mathbb{F}_{2^{n}}^{2}\\ u\neq 0\end{subarray}}(-1)^{\operatorname{Tr}(u)+\operatorname{Tr}(vb)}W_{F}^{3}(1,u^{-d}a^{d}v)+\sum_{v\in\mathbb{F}_{2^{n}}}(-1)^{\operatorname{Tr}(vb)}W_{F}^{3}(0,v)$
$=\sum_{\begin{subarray}{c}(u,v)\in\mathbb{F}_{2^{n}}^{2}\\ u\neq 0\end{subarray}}(-1)^{\operatorname{Tr}(u)+\operatorname{Tr}(va^{-d}b)}W_{F}^{3}(1,u^{-d}v)+\sum_{v\in\mathbb{F}_{2^{n}}}(-1)^{\operatorname{Tr}(vb)}W_{F}^{3}(0,v)$
$=\sum_{\begin{subarray}{c}(u,v)\in\mathbb{F}_{2^{n}}^{2}\\ u\neq 0\end{subarray}}(-1)^{\operatorname{Tr}(u)+\operatorname{Tr}(va^{-d}b)}W_{F}^{3}(u,v)+\sum_{v\in\mathbb{F}_{2^{n}}}(-1)^{\operatorname{Tr}(vb)}W_{F}^{3}(0,v).$
We then have that
$\widehat{W_{F}^{3}}(a,b)=\widehat{W_{F}^{3}}(1,a^{-d}b)-\sum_{v\in\mathbb{F}_{2^{n}}}(-1)^{\operatorname{Tr}(va^{-d}b)}W_{F}^{3}(0,v)+\sum_{v\in\mathbb{F}_{2^{n}}}(-1)^{\operatorname{Tr}(vb)}W_{F}^{3}(0,v).$
However, notice that the following is equal to 0
$-\sum_{v\in\mathbb{F}_{2^{n}}}(-1)^{\operatorname{Tr}(va^{-d}b)}W_{F}^{3}(0,v)+\sum_{v\in\mathbb{F}_{2^{n}}}(-1)^{\operatorname{Tr}(vb)}W_{F}^{3}(0,v)$
$=\sum_{v\in\mathbb{F}_{2^{n}}}(-1)^{\operatorname{Tr}(vb)}\left(W_{F}^{3}(0,v)-W_{F}^{3}(0,a^{d}v)\right)$
by the equality $W_{F}(0,a^{d}v)=W_{F}(0,v)$. Therefore $\widehat{W_{F}^{3}}(a,b)=\widehat{W_{F}^{3}}(1,a^{-d}b)$, implying $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\operatorname{mult}_{\mathcal{G}_{F}}(1,a^{-d}b)$ as $\operatorname{mult}_{\mathcal{G}_{F}}(x,y)=\frac{1}{6\cdot 2^{2n}}\widehat{W_{F}^{3}}(x,y)$ for all $(x,y)\notin\mathcal{G}_{F}$. Note that $b\mapsto a^{-d}b$ is a permutation, and so the exclude multiplicities contained in $\left\{a\right\}\times(\mathbb{F}_{2^{n}}\setminus\left\{F(a)\right\}$ are exactly those that are contained in $\left\{a\right\}\times(\mathbb{F}_{2^{n}}\setminus\left\{1\right\})$. Finally, we can apply our equality twice to see that for any $c\neq 0$, we have $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\operatorname{mult}_{\mathcal{G}_{F}}(1,a^{-d}b)=\operatorname{mult}_{\mathcal{G}_{F}}(c,(a^{-1}c)^{d}b)$. ∎
If $F(x)=x^{d}$ is a APN and $n$ is odd, we can determine the exact exclude multiplicities of $\mathcal{G}_{F}$ contained in $\left\{0\right\}\times\mathbb{F}_{2^{n}}^{\ast}$, but first, we describe the sum $\sum_{x\in\mathbb{F}_{2^{n}}}\gamma_{F}(x+a,x^{d}+b)$ in the following way to make our computation straightforward.
Proposition 5.2.
Let $F\colon\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ be a power function $F(x)=x^{d}$, and let $(a,b)\in\mathbb{F}_{2^{n}}$. Let $\Phi\colon\mathbb{F}_{2^{n}}\setminus\left\{a\right\}\to\mathbb{F}_{2^{n}}$ be the function $\Phi(x)=\frac{x^{d}+b}{(x+a)^{d}}$. Then
$\sum_{x\in\mathbb{F}_{2^{n}}}\gamma_{F}(y+a,y^{d}+b)=|\Phi^{-1}(\operatorname{Im}D_{1}F)|.$
Proof.
When $y\neq a$, we have $D_{y+a}F(x)=(y+a)^{d}D_{1}F\left(\frac{x}{y+a}\right)$. Therefore if $y\neq a$, then $\gamma_{F}(y+a,y^{d}+b)=1$ if and only if $D_{1}F\left(\frac{x}{y+a}\right)=\frac{y^{d}+b}{(y+a)^{d}}$ has a solution. For any fixed $y\neq a$, the function $x\mapsto\frac{x}{y+a}$ is a permutation. Therefore the sum $\sum_{x\in\mathbb{F}_{2^{n}}}\gamma_{F}(y+a,y^{d}+b)$ is the number of $y\neq a$ such that $\frac{y^{d}+b}{(y+a)^{d}}$ is contained in $\operatorname{Im}D_{1}F$, that is, the size of the preimage of $\operatorname{Im}D_{1}F$ under $\Phi$. Hence
$\sum_{x\in\mathbb{F}_{2^{n}}}\gamma_{F}(y+a,y^{d}+b)=|\Phi^{-1}(\operatorname{Im}D_{1}F)|,$
as desired. ∎
With this fact and our previous results, we will be able to determine the exclude multiplicities of $\mathcal{G}_{F}$ that are contained in $\left\{0\right\}\times\mathbb{F}_{2^{n}}^{\ast}$ when $n$ is odd. This is because in Proposition 3.3, we proved that $3\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\sum_{x\in\mathbb{F}_{2}^{n}}\gamma_{F}(x+a,F(x)+b)$ for all $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$.
Proposition 5.3.
Assume $n\geq 3$. Let $F\colon\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ be an APN power function $F(x)=x^{d}$. If $n$ is odd, then for all $a\neq 0$, we have
$\operatorname{mult}_{\mathcal{G}_{F}}(a,0)=\operatorname{mult}_{\mathcal{G}_{F}}(0,a)=\frac{2^{n-1}-1}{3}.$
Proof.
Assume $n$ is odd. For $(a,b)\in\mathbb{F}_{2^{n}}^{2}$, let $\Phi\colon\mathbb{F}_{2^{n}}\setminus\left\{a\right\}\to\mathbb{F}_{2^{n}}$ be defined by $\Phi(x)=\frac{x^{d}+b}{(x+a)^{d}}$. First, suppose $a\neq 0$ and $b=0$. Then $\Phi(x)=\frac{x^{d}}{(x+a)^{d}}=\left(\frac{x}{x+a}\right)^{d}$. For $x\neq a$, we have $\frac{x}{x+a}=a(x+a)^{-1}+1$. Note $x\mapsto a(x+a)^{-1}+1$ is a permutation from $\mathbb{F}_{2^{n}}\setminus\left\{a\right\}$ is a permutation onto its image which is $\mathbb{F}_{2^{n}}\setminus\left\{1\right\}$. So $|\Phi^{-1}(\operatorname{Im}D_{1}F)|=|\left\{x\in\mathbb{F}_{2^{n}}\setminus\left\{1\right\}:x^{d}\in\operatorname{Im}D_{1}F\right\}|$. Note that $F(x)=x^{d}$ is a permutation since $n$ is odd. Hence $1\notin\operatorname{Im}\Phi$ but $1\in\operatorname{Im}D_{1}F$. Applying Proposition 3.3 and Proposition 5.2, we have
$\operatorname{mult}_{\mathcal{G}_{F}}(a,0)=\frac{|\Phi^{-1}(\operatorname{Im}D_{1}F)|}{3}=\frac{|\operatorname{Im}D_{1}F|-1}{3}=\frac{2^{n-1}-1}{3}.$
Now, suppose $a=0$ and $b\neq 0$. Then $\Phi(x)=\frac{x^{d}+b}{x^{d}}=bx^{-d}+1$. Hence $\Phi$ is a permutation onto its image which is $\mathbb{F}_{2^{n}}\setminus\left\{1\right\}$, and similar to before we have $1\notin\operatorname{Im}\Phi$ but $1\in\operatorname{Im}D_{1}F$. Therefore
$\operatorname{mult}_{\mathcal{G}_{F}}(0,a)=\frac{|\Phi^{-1}(\operatorname{Im}D_{1}F)|}{3}=\frac{2^{n-1}-1}{3}.$
In [27, Conjecture 21], Kaleyski conjectured that if $F$ is an APN power function, the two values $\sum_{y\in\mathbb{F}_{2^{n}}}\delta_{F}(y,F(y)+1)$ and $\sum_{y\in\mathbb{F}_{2^{n}}}\delta_{F}(y+1,F(y))$ are equal to $\sum_{y\in\mathbb{F}_{2^{n}}}\delta_{x^{3}}(y,y^{3}+1)$. By Proposition 3.3, this is equivalent to $\operatorname{mult}_{\mathcal{G}_{F}}(0,1)=\operatorname{mult}_{\mathcal{G}_{F}}(1,0)=\operatorname{mult}_{\mathcal{G}_{x^{3}}}(0,1)$. Therefore, Proposition 5.3 proves Kaleyski’s conjecture in the case of $n$ odd, but the case of $n$ even still remains open.
Remark 5.4.
The reader may have noticed that the particular case of $\Phi(x)=\frac{x^{d}+b}{(x+a)^{d}}$ when $a=b=1$ was studied by Carlet and Picek in [15]. In particular, they proved $x\mapsto x^{d}$ is APN if and only if $x\mapsto\frac{x^{d}+1}{(x+1)^{d}}$ is 22-to-11 from $\mathbb{F}_{2^{n}}\setminus\mathbb{F}_{2}\to\mathbb{F}_{2^{n}}\setminus\left\{1\right\}$. More generally, it is true that $F(x)=x^{d}$ is APN if and only if for all $a\in\mathbb{F}_{2^{n}}^{\ast}$ the map $\Phi(x)=\frac{x^{d}+a^{d}}{(x+a)^{d}}$ is 22-to-11 from $\mathbb{F}_{2^{n}}\setminus\left\{0,a\right\}$ to $\mathbb{F}_{2^{n}}\setminus\left\{1\right\}$. To see this, suppose $x\neq a$ and let $s=\frac{x}{x+a}$. Observe that $x\mapsto s$ is a permutation from $\mathbb{F}_{2^{n}}\setminus\left\{a\right\}$ to $\mathbb{F}_{2^{n}}\setminus\left\{1\right\}$. Also $D_{1}F(s)=s^{d}+(s+1)^{d}=\frac{x^{d}+a^{d}}{(x+a)^{d}}$. Then $x\mapsto\frac{x^{d}+a^{d}}{(x+a)^{d}}$ is 22-to-11 from $\mathbb{F}_{2^{n}}\setminus\left\{0,a\right\}$ to $\mathbb{F}_{2^{n}}\setminus\left\{a\right\}$ if and only if for any $b\neq 1$, the equation $D_{1}(s)=b$ has at most $2$ solutions $s\in\mathbb{F}_{2^{n}}$ (and there are no solutions in $\left\{0,a\right\}$). Therefore $x\mapsto\frac{x^{d}+a^{d}}{(x+a)^{d}}$ is 22-to-11 from $\mathbb{F}_{2^{n}}\setminus\left\{0,a\right\}$ to $\mathbb{F}_{2^{n}}\setminus\left\{a\right\}$ if and only if $D_{1}F$ is 22-to-11, which is equivalent to $F$ being APN.
6 Lower bounds on distances from the APN inverse function
Throughout this section, we assume that $n$ is odd and define $F\colon\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ to be the multiplicative function $F(x)=x^{-1}$, where $\frac{1}{0}:=0$. To determine lower bounds on the Hamming distance from $F(x)=x^{-1}$ over $\mathbb{F}_{2^{n}}$ to any other APN function, we will determine lower bounds on the multiplicities of the exclude points of $\mathcal{G}_{F}$. In particular, we derive a lower bound in terms of the binary Kloosterman sums
$K_{n}(a)=\sum_{x\in\mathbb{F}_{2^{n}}}(-1)^{\operatorname{Tr}(ax+x^{-1})}.$
For notational convenience, we also denote by $K_{n}^{\ast}(a)$ the sum
$K_{n}^{\ast}(a)=\sum_{x\in\mathbb{F}_{2^{n}}^{\ast}}(-1)^{\operatorname{Tr}(ax+x^{-1})}.$
A well-known result is that $\left\{K_{n}^{\ast}(a):a\neq 0\right\}$ is the set of integers in $2^{\frac{n}{2}+1},2^{\frac{n}{2}+1}]$ that are congruent to $3\mod 4$, see [32]. Kloosterman sums and the Walsh transform of $F$ have a very close connection as one can easily verify that $W_{F}(u,v)=K_{n}(uv)+2^{n}\delta_{(0,0)}(u,v)$.
Kloosterman sums can also be used to describe the number of rational points on ordinary elliptic curves over $\mathbb{F}_{2^{n}}$. First, recall that an algebraic plane curve $\mathcal{C}$ over $\mathbb{F}_{2^{n}}$ is defined by the equation $p(x,y)=0$ for some irreducible polynomial $p(x,y)$ over $\mathbb{F}_{2^{n}}$. If $\widetilde{p}(x,y,z)$ is the associated homogenized polynomial to $p(x,y)$ we denote the solution set to $\widetilde{p}(x,y,z)=0$ as $\widetilde{C}$. We say that the number of rational points of $C$ is the size of $\widetilde{C}$. Also recall that a point $(x,y,z)\in\widetilde{C}$ is singular if all partial derivatives of $\widetilde{p}(x,y,z)$ vanish at $(x,y,z)$. If there are no singular points, we define the genus of $\mathcal{C}$ as $g=\frac{(d-1)(d-2)}{2}$, where $d$ is the degree of $\widetilde{p}(x,y,z)$. If the genus of $\mathcal{C}$ is $1$, then we say that $\mathcal{C}$ is elliptic. Moreover, we say that $\mathcal{C}$ is supersingular if the coefficient of xyzxyz is zero in $\widetilde{p}(x,y,z)$, otherwise we say that $\mathcal{C}$ is ordinary (cf. [32]).
Lachaud and Wolfmann proved the following two results in [32].
Proposition 6.1 ([32][Corollary 2.2]).
An ordinary elliptic curve $\mathcal{E}$ defined over $\mathbb{F}_{2^{n}}$ is isomorphic to one of the following Kloosterman curves
$\mathcal{KL}^{+}_{a}$ $\colon y^{2}+y=ax+x^{-1}\quad a\in\mathbb{F}_{2^{n}}^{\ast},$
$\mathcal{KL}^{-}_{a}$ $\colon y^{2}+y=ax+x^{-1}+\tau\quad a,\tau\in\mathbb{F}_{2^{n}},a\neq 0.$
Proposition 6.2 ([32][Corollary 3.3]).
For any $a\neq 0$, the number of rational points of $\mathcal{KL}^{\pm}_{a}$ is given by $2^{n}+1\pm K_{n}^{\ast}(a)$.
We will now prove the following theorem and then use Lemma 3.1 to obtain lower bounds on the distance of any other APN function to $F$.
Theorem 6.3.
Assume $n\geq 3$ is odd. Let $F\colon\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ be the function $F(x)=x^{-1}$ where we define $F(0)=0$. Then for all $(a,b)\in\mathbb{F}_{2^{n}}^{2}\setminus\mathcal{G}_{F}$, we have
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq\left\lceil\frac{2^{n}-5-\max_{a\in\mathbb{F}_{2^{n}}^{\ast}}|K_{n}(a)-1|}{6}\right\rceil$ (10)
Proof.
Let $(a,b)\notin\mathcal{G}_{F}.$ The exclude multiplicity of $(a,b)$ is $\frac{s}{6}$ where $s$ counts the number of solutions $(x,y,z)\in\mathbb{F}_{2^{n}}^{3}$ to
$\begin{cases}x+y+z=a\\ x^{-1}+y^{-1}+z^{-1}=b,\end{cases}$
Equivalently, $s$ is the number of roots $(x,y)\in$ $\mathbb{F}_{2^{n}}^{2}$ of the polynomial
$p(x,y)=x^{-1}+y^{-1}+(x+y+a)^{-1}+b.$ (11)
By Proposition 5.3, we know that if either $a$ or $b$ is zero, then $s=2^{n}-2$ since $F$ is APN and $n$ is odd. Moreover, as shown in Proposition 5.1, the values of $s$ are independent of $a$ as $a$ ranges across $\mathbb{F}_{2^{n}}^{\ast}$. So, without loss of generality, assume $a=1$ and $b\notin\mathbb{F}_{2}$. Define the affine plane curve $\mathcal{C}=\left\{(x,y)\in\mathbb{F}_{2^{n}}^{2}:p(x,y)=0\right\},$ and let $D\subset\mathcal{C}$ be the subset where no denominator vanishes:
$D=\left\{(x,y)\in\mathcal{C}:0\notin\left\{x,y,x+y+1\right\}\right\}$
We claim that $\mathcal{C}=|D|+3\delta_{F}(1,b)$. Indeed, if $x=0$, then $p(0,y)=y^{-1}+(y+1)^{-1}+b=0$ has $\delta_{F}(1,b)$ solutions. Similarly, the cases $y=0$ and $x+y=1$ each contribute $\delta_{F}(1,b)$ distinct solutions. Hence, $\mathcal{C}=|D|+3\delta_{F}(1,b)$.
Now we multiply eq. 11 by $xy(x+y+1)$ to obtain the following polynomial in $\mathbb{F}_{2^{n}}[x,y]$:
$q(x,y)=bx^{2}y+bxy^{2}+x^{2}+y^{2}+bxy+xy+x+y.$ (12)
Let $\mathcal{E}=\{(x,y)\in\mathbb{F}_{2^{n}}^{2}:q(x,y)=0\}$, and let $G=\left\{(x,y)\in E:0\notin\left\{x,y,x+y+1\right\}\right\}$. Note that for all $(x,y)\in\mathbb{F}_{2^{n}}^{2}$ such that $x,y,x+y+1\neq 0$, we have $p(x,y)=0$ if and only if $q(x,y)=0.$ Hence $|D|=|G|$.
We claim that $\mathcal{E}=|G|+3$. To see this, let us consider a point $(x,y)\in\mathcal{E}\setminus G$. If $x=0,$ then $0=q(0,y)=y^{2}+y.$ Hence $y\in\mathbb{F}_{2}$. Similarly, if $y=0,$ then $x\in\mathbb{F}_{2}$. Finally, if $x+y=1,$ then
$0=bx^{2}(x+1)+bx(x+1)^{2}+bx(x+1)+x(x+1)+x^{2}+(x+1)^{2}+x+(x+1)=x^{2}+x.$
This implies $x\in\mathbb{F}_{2}$, and similarly $y\in\mathbb{F}_{2}$. Since $q(1,1)=b+1\neq 0$, we have $(x,y)\in\mathbb{F}_{2}^{2}\setminus\left\{(1,1)\right\}$. Therefore $\mathcal{E}=|G|+3$.
Now, we homogenize $q(x,y)$ to obtain the homogeneous polynomial
$\tilde{q}(x,y,z)=bx^{2}y+bxy^{2}+x^{2}z+y^{2}z+(b+1)xyz+xz^{2}+yz^{2},$
Let $\widetilde{\mathcal{E}}=\{(x,y,z):\tilde{q}(x,y,z)=0\}$ denote the projective curve defined by $\tilde{q}$. By considering the partial derivatives of $\tilde{q}$, it is easy to see that $\mathcal{E}$ is smooth, and we know that $\mathcal{E}$ is elliptic because its genus is $g=\frac{(3-1)(3-2)}{2}=1$. Moreover, $\mathcal{E}$ is ordinary because xyzxyz has a nonzero coefficient in $\tilde{q}$. Note that the projective points of $\widetilde{\mathcal{E}}$ include three additional points, corresponding to solutions of the form $(x:y:0)$ that satisfy $\widetilde{q}(x,y,0)=bx^{2}y+bxy^{2}=bxy(x+y)=0$. There are three such cases: $x=0,y=0$ and $x+y=0,$ corresponding to the points $(1:0:0),(0:1:0),(1:1:0)\in\widetilde{E}$ respectively. So $\widetilde{\mathcal{E}}=\#\mathcal{E}+3$.
Hence, $6\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=s=\#\mathcal{C}=|D|+3\delta_{F}(1,b)=|G|+3\delta_{F}(1,b)=\#\mathcal{E}-3+3\delta_{F}(1,b)$, implying $6\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\#\widetilde{\mathcal{E}}-6+3\delta_{F}(1,b)$. By Proposition 6.2, we have that $\widetilde{\mathcal{E}}$ is equal to $2^{n}+1\pm K_{n}^{\ast}(u)$ for some $u\in\mathbb{F}_{2^{n}}^{\ast}$ as $\widetilde{\mathcal{E}}$ is an ordinary elliptic curve. So $\widetilde{\mathcal{E}}\geq\min_{v\in\mathbb{F}_{2^{n}}^{\ast}}\left(2^{n}+1\pm K_{n}^{\ast}(v)\right)=2^{n}+1-\max_{v\in\mathbb{F}_{2^{n}}^{\ast}}|K_{n}^{\ast}(v)|$. Therefore,
$6\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ $\geq 2^{n}-5-\max_{a\in\mathbb{F}_{2^{n}}^{\ast}}|K_{n}^{\ast}(a)|+3\delta_{F}(1,b)$
$\geq 2^{n}-5-\max_{a\in\mathbb{F}_{2^{n}}^{\ast}}|K_{n}(a)-1|$
and eq. 10 immediately follows. ∎
Therefore, we have obtained lower bounds on the exclude multiplicity of any point in $\mathbb{F}_{2^{n}}^{2}\setminus\mathcal{G}_{F}$, with our lower bound being approximately $\frac{2^{n}-2^{\frac{n}{2}+1}}{6}$. Moreover, we provide explicit computations of the exact minimal exclude multiplicity of $\mathcal{G}_{F}$ for $3\leq n\leq 15$ (with $n$ odd) compared to our lower bound in Table 4.
$n$ $e_{\min}(\mathcal{G}_{F})$ $\left\lceil\frac{1}{6}\left(2^{n}-5-\max_{a\in\mathbb{F}_{2^{n}}^{\ast}}|K_{n}(a)-1|\right)\right\rceil$
3 1 0
5 3 3
7 18 17
9 77 77
11 326 326
13 1335 1335
15 5401 5401
Table 4: The minimum exclude multiplicities of the graph of the APN inverse function compared to the lower bound of Theorem 6.3 for odd $3\leq n\leq 15$.
Proof of theorem 1.3.
Apply Lemma 3.1 to Theorem 6.3. ∎
7 On the existence of linear structures of $\gamma_{F}$
For $\epsilon\in\mathbb{F}_{2}$, we say that an $\epsilon$-valued linear structure of a Boolean function $f\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ is a nonzero point $a\in\mathbb{F}_{2}^{n}$ such that $D_{a}f(x)=f(x)+f(x+a)$ takes constant value $\epsilon$. Since $\gamma_{F}$ has weight $\binom{2^{n}}{2}$ when $F$ is APN (see [14]), it is not balanced, and therefore cannot have a $1$-valued linear structure. However, not very much is known about whether or not $\gamma_{F}$ can have 0-valued linear structures.
In [13], there were two classes of APN functions that were shown that do not admit linear structures: the first if when $F$ is an APN power function, and the second is when $F$ is such that there exists some $x\in\mathbb{F}_{2}^{n}$ with the property that $\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:D_{a}D_{b}F(x)=w\right\}$ does not depend on $w\in\mathbb{F}_{2}^{n}\setminus\left\{0\right\}$.
We will prove that if $F$ is a plateaued 3-to-1 function ($n$ even), then $\gamma_{F}$ does not admit nontrivial linear structures, forming the third known class of APN functions such that $\gamma_{F}$ has no nontrivial linear structures. Let us rephrase the problem of $\gamma_{F}$ having linear structures in terms of exclude multiplicity when $F$ is a plateaued APN function.
Proposition 7.1.
Assume $n\geq 4$ is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be plateaued APN function. Then $\gamma_{F}$ has a nontrivial linear structure if and only if there exists $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$ such that $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\lfloor\frac{2^{n}}{3}\rfloor=\frac{2^{n}-1}{3}$. Equivalently, $\gamma_{F}$ has a nontrivial linear structure if and only if $\mathcal{G}_{F}$ has at least $2^{n}$ points of the maximum possible exclude multiplicity.
Proof.
Let $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$. By Proposition 4.17, we know that
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\frac{2^{n}-2}{6}+\frac{2^{2n}-2\operatorname{wt}(D_{(0,b+F(a))}\gamma_{F})}{6\cdot 2^{n}}$
Therefore $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\frac{2^{n}-1}{3}$ if and only if $\operatorname{wt}(D_{(0,b+F(a))}\gamma_{F})=0$. ∎
Proposition 7.1 shows that obtaining an upper bound on the exclude multiplicities of the graph of a plateaued APN function also poses as a difficult problem. However, recall that in Section 4.3, we determined the exact exclude multiplicities that occur when $F$ is a plateaued 3-to-1 function. The following result is then an immediate corollary.
Proposition 7.2.
Assume $n\geq 4$ and is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued 3-to-1 function. Then $\gamma_{F}$ has no nontrivial linear structures.
Proof.
By Theorem 4.20, any exclude point of $\mathcal{G}_{F}$ has multiplicity $\frac{2^{n}+(-2)^{\frac{n}{2}+1}-2}{6}$ or $\frac{2^{n}+(-2)^{\frac{n}{2}}-2}{6}$, and so any exclude point of $\mathcal{G}_{F}$ has multiplicity strictly less than $\frac{2^{n}-1}{3}$. The result then follows from Proposition 7.1. ∎
As shown above, there is an equivalence of $\gamma_{F}$ having a nontrivial linear structure and $\mathcal{G}_{F}$ having an exclude point of multiplicity $\frac{2^{n}-1}{3}$ when $F$ is plateaued APN. In the case of $F$ also being quadratic, we can then characterize the existence of a nontrivial linear structure of $\gamma_{F}$ in terms of the components of the ortho-derivative of $F$.
Corollary 7.3.
Suppose $n\geq 4$ is even, and let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a quadratic APN function. Then $\gamma_{F}$ has a nontrivial linear structure if and only if the ortho-derivative of $F$ has a component function of weight 0. Equivalently, $\gamma_{F}$ has a nontrivial linear structure if and only if the image of $\pi_{F}$ is contained in a linear hyperplane.
Proof.
Apply Proposition 7.1 to eq. 9. ∎
In particular, the class of quadratic APN functions whose ortho-derivatives have nonzero nonlinearity is a subclass of those APN functions such that $\gamma_{F}$ has no nontrivial linear structures. This immediately follows by applying the inequality $\operatorname{wt}(b\cdot F)\geq\operatorname{\mathcal{NL}}(\pi_{F})$ for $b\neq 0$ to eq. 9. In particular, we have the inequality
$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\leq\frac{2^{n}-1-\operatorname{\mathcal{NL}}(\pi_{F})}{3}.$ (13)
when $F$ is quadratic APN and $(a,b)\notin\mathcal{G}_{F}$. By combining eq. 13 and Proposition 4.17, we have that for all even $n$ and any quadratic APN function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ such that $F(0)=0$ and $b\neq 0$ such, we have $\Delta_{\gamma_{F}}(0,b)\leq 2^{2n}-2^{n+1}\operatorname{\mathcal{NL}}(\pi_{F})$, or equivalently
$\operatorname{wt}(D_{(0,b)}\gamma_{F})\geq 2^{n}\operatorname{\mathcal{NL}}(\pi_{F}).$
Also, by Theorem 4.20, one can easily deduce the exact weights of $D_{(0,b)}\gamma_{F}$ for all $b\neq 0$ when $F$ is a plateaued 33-to-11 function. Moreover, we have the following inequalities on the weights of the derivatives of $\gamma_{F}$ of the form $D_{(0,b)}\gamma_{F}$.
Proposition 7.4.
Suppose $n\geq 4$ is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a quadratic APN function with $F(0)=0$. Then for any $b\neq 0$, the weight of $D_{(0,b)}\gamma_{F}$ is divisible by $6\cdot 2^{n}$ and satisfies
$0\leq\operatorname{wt}(D_{(0,b)}\gamma_{F})\leq 2^{n}(2^{n}-3\cdot 2^{\frac{n}{2}-1}+2).$
Proof.
By Proposition 4.25, we know $\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=2k+1$ for some non-negative integer $k\geq 0$. Therefore, by Proposition 4.17, we have $6\cdot 2^{n}(2k+1)=2^{n}(2^{n}-2)+2^{2n}-2\operatorname{wt}(D_{(0,b)}\gamma_{F})$. So $\operatorname{wt}(D_{(0,b)}\gamma_{F})=2^{n}(2^{n}-6k-4)$. Since $2^{n}-4$ is divisible by $6$ as $n$ is even, we have $\operatorname{wt}(D_{(0,b)}\gamma_{F})\equiv 0\mod 6\cdot 2^{n}$.
By Theorem 4.13, we have $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq 2^{\frac{n}{2}-1}-1$ for all $(a,b)\notin\mathcal{G}_{F}$ as $F$ is quadratic and therefore plateaued. Moreover, we have $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\leq\frac{2^{n}-1}{3}$ by [19, Corollary 3.4]. So, for $b\neq 0$, we have $2^{\frac{n}{2}-1}-1\leq\operatorname{mult}_{\mathcal{G}_{F}}(0,b)\leq\frac{2^{n}-1}{3}$, and the result follows. ∎
8 Open problems
To conclude the paper, we list a few open problems relating to the exclude multiplicities of the graphs of APN functions.
Open Problem 8.1.
Establish a lower bound on the exclude multiplicities of the graph of the Dobbertin function or all APN power functions in general.
Open Problem 8.2.
Establish a better lower bound on the exclude multiplicities of the graph of a plateaued (or quadratic) APN function or find an infinite family that tightly meets our bound from Theorem 4.13.
Open Problem 8.3.
For $n$ even, prove or disprove the existence of a plateaued (or quadratic) APN functions whose graph has an exclude point of multiplicity $\frac{2^{n}-1}{3}$ (the maximum that is theoretically possible). Recall that Proposition 7.1 characterizes this problem in terms of the existence of a nontrivial linear structure of $\gamma_{F}$.
Open Problem 8.4.
Study the nonlinearity of the ortho-derivative $\pi_{F}$ of a quadratic APN function $F$. Is it possible for $\pi_{F}$ to have zero nonlinearity?
Open Problem 8.5.
For $n$ even, find an example of a non-quadratic APN function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ whose graph has an exclude point of even multiplicity, or prove such a function does not exist. Note that if such a function does not exist, then 2.1 follows for all even $n$. It may be easier to first prove or disprove 4.15.
Acknowledgment
We thank Claude Carlet and Larry Rolen for useful discussions and helpful feedback that improved earlier drafts of this paper.

### v2

@@
 Lower bounds on the distance from an APN function $F$ to any other are known for almost bent (AB) functions and when $F$ is a 3-to-1 quadratic function with $n$ even.
-In this paper, we reinterpret $\Pi_{F}$ in terms of the exclude multiplicities of the graph $\mathcal{G}_{F}=\left\{(x,F(x)):x\in\mathbb{F}_{2}^{n}\right\}$ of $F$ as a Sidon set.
+In this paper, we reinterpret $\Pi_{F}$ in terms of the exclude multiplicities of the graph $\mathcal{G}_{F}=\{(x,F(x)):x\in\mathbb{F}_{2}^{n}\}$ of $F$ as a Sidon set.
 We establish lower bounds on the distance for even $n$ when $F$ is plateaued APN, generalize the known lower bounds for quadratic 33-to-11 function to all 3-to-1 plateaued functions (e.g.
@@
 The proofs of each of our main results all rely on finding a lower bound on $e_{\min}(\mathcal{G}_{F})$, and in some cases, we determine the exact exclude multiplicities of $\mathcal{G}_{F}$.
+The lower bounds on the distances for different families of APN functions are summarized in Table 1.
 $F\colon\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ Lower bounds of $d(F,G)$ for $G\neq F$ and $G$ APN Reference
@@
 As a corollary, we prove that the Brinkmann-Leander-Edel-Pott function (the only known instance of an APN function not CCZ-equivalent to a power function or a quadratic function) is also not CCZ-equivalent to a plateaued function.
+We also prove that the exclude multiplicities of $\mathcal{G}_{F}$ are odd when $F$ is plateaued APN ($n\geq 3$) via a characterization in terms of the intersections of ${\mathcal{B}({F})}=\left\{v\in\mathbb{F}_{2}^{n}:v\cdot F\text{ is bent}\right\}$ with affine hyperplanes.
 In Section 4.4, we study quadratic APN functions with a focus on how their ortho-derivatives can be used to express exclude multiplicities.
-By using a recently proved fact about the algebraic degree of an ortho-derivative, we show that all exclude multiplicities of $\mathcal{G}_{F}$ are odd when $F$ is quadratic APN.
-As an equivalent result, we prove that ${\mathcal{B}({F})}=\left\{v\in\mathbb{F}_{2}^{n}:v\cdot F\text{ is bent}\right\}$ always has an intersection of even size with any affine hyperplane of $\mathbb{F}_{2}^{n}$ when $F$ is quadratic APN.
 In Section 5, we study the exclude multiplicities of APN power functions, and we show that $\Pi_{F}$ is completely determined by $\Pi_{F}^{0}$ and $\Pi_{F}^{1}$.
@@
 In Section 7, we study how exclude multiplicity is related to the $\gamma_{F}$ Boolean function associated to an APN function $F$.
-In particular, we show that if $F$ is plateaued APN, then $\gamma_{F}$ has no nontrivial linear structures if and only if $\mathcal{G}_{F}$ does not have a point of exclude multiplicity $\frac{2^{n}-1}{3}$ (the largest possible).
-As a consequence, we prove that $\gamma_{F}$ has no nontrivial linear structures when $F$ is plateaued 3-to-1.
-We also show that $\gamma_{F}$ has no nontrivial linear structures when $F$ is quadratic APN with an ortho-derivative of nonzero nonlinearity.
+In particular, we show that if $F$ is plateaued APN, then $\gamma_{F}$ has no nontrivial linear structures if and only if $\mathcal{G}_{F}$ does not have a point of exclude multiplicity $\frac{2^{n}-1}{3}$ (the largest possible), and we show that this latter condition always holds, proving that $\gamma_{F}$ has no nontrivial linear structures when $F$ is plateaued APN.
 To conclude the paper, we list open problems in Section 8.
@@
 Recall that the Fourier-Hadamard transform of a function $\varphi\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}$ is the function $\widehat{\varphi}(u)=\sum_{x\in\mathbb{F}_{2}^{n}}(-1)^{x\cdot u}\varphi(x)$.
-We will also use the convolutional product, which for any two functions $\varphi,\psi\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}$ is given by $(\varphi\times\psi)(u)=\sum_{x\in\mathbb{F}_{2}^{n}}\varphi(x)\psi(x+u)$.
+We will also use the convolutional product, which for any two functions $\varphi,\psi\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}$ is given by $(\varphi\otimes\psi)(u)=\sum_{x\in\mathbb{F}_{2}^{n}}\varphi(x)\psi(x+u)$.
 For a Boolean function $f\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$, we say that its (Hamming) weight is the size of its support, i.e.
@@
 A Sidon set is called maximal if it is not contained in any strictly larger Sidon set.
-In [12] used the connection of APN functions and Sidon sets to prove that 2.1 is equivalent to the following conjecture.
+In [16], Carlet used the connection of APN functions and Sidon sets to prove that 2.1 is equivalent to the following conjecture.
 Conjecture 2.2 ([16]).
@@
 We now compute the size of $\Pi_{F}^{\beta}(b)$ in terms of $\gamma_{F}$, and this will be useful in demonstrating the connection between shifted derivatives and exclude multiplicity.
-Note that for any $u\in\mathbb{F}_{2}^{n}$, we write the Kronecker delta as $\delta_{0}(u)$, which takes value $1$ if and only if $u=0$ and value 0 otherwise.
+Note that for any $u\in\mathbb{F}_{2}^{n}$, we write the Dirac delta as $\delta_{0}(u)$, which takes value $1$ if and only if $u=0$ and value 0 otherwise.
 Proposition 3.3.
@@
 $=\frac{1}{2}((1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})\otimes 1_{\mathcal{G}_{F}})(\beta,b)-2^{n-1}1_{\mathcal{G}_{F}}(\beta,b)$
-Since $\widehat{\varphi\otimes\psi}=\widehat{\varphi}\widehat{\psi}$ for any functions $\varphi,\psi\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}$ (see [11, Proposition 11]) and using the fact that $\widehat{1_{\mathcal{G}_{F}}}=W_{F}$, the Fourier-Hadamard transform of $(1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})\otimes 1_{\mathcal{G}_{F}}$ is $\widehat{(1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})}\widehat{1_{\mathcal{G}_{F}}}=(\widehat{1_{\mathcal{G}_{F}}}\widehat{1_{\mathcal{G}_{F}}})\widehat{1_{\mathcal{G}_{F}}}=W_{F}^{3}$.
+Since $\widehat{\varphi\otimes\psi}=\widehat{\varphi}\widehat{\psi}$ for any functions $\varphi,\psi\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}$ (see [14, Proposition 11]) and using the fact that $\widehat{1_{\mathcal{G}_{F}}}=W_{F}$, the Fourier-Hadamard transform of $(1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})\otimes 1_{\mathcal{G}_{F}}$ is $\widehat{(1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})}\widehat{1_{\mathcal{G}_{F}}}=(\widehat{1_{\mathcal{G}_{F}}}\widehat{1_{\mathcal{G}_{F}}})\widehat{1_{\mathcal{G}_{F}}}=W_{F}^{3}$.
 Therefore, taking the Fourier-Hadamard transform of both sides, we have $(1_{\mathcal{G}_{F}}\otimes 1_{\mathcal{G}_{F}})\otimes 1_{\mathcal{G}_{F}}=\frac{1}{2^{2n}}\widehat{W_{F}^{3}}$, and so
@@
 $=\frac{1}{2^{2n+1}}\widehat{W_{F}^{3}}(\beta,b)-\delta_{0}(b+F(\beta))(2^{n-1}-1).$
-In particular, when $(\beta,b)\notin\mathcal{G}_{F}$ and $F$ is APN, we have
+In particular, when $(\beta,b)\notin\mathcal{G}_{F}$, we have
 $|\Pi_{F}^{\beta}(b)|=\frac{1}{2^{2n+1}}\widehat{W_{F}^{3}}(\beta,b)=3\operatorname{mult}_{\mathcal{G}_{F}}(\beta,b).$
@@
 $d(F,G)\geq\left\lceil\frac{m_{F}}{3}\right\rceil+1.$
-For any APN function $F$, any exclude point of $\mathcal{G}_{F}$ has exclude multiplicity at most $\lfloor\frac{2^{n}}{3}\rfloor$ (see [19, Corollary 3.4]).
+For any APN function $F$, any exclude point of $\mathcal{G}_{F}$ has exclude multiplicity at most $\lfloor\frac{2^{n}}{3}\rfloor$ (see [20, Corollary 3.4]).
 Therefore, $m_{F}=\min\left\{3e_{\min}(\mathcal{G}_{F}),2^{n}\right\}=3e_{\min}(\mathcal{G}_{F})$, and so the bound of Corollary 3.5 is equivalent to the bound of Lemma 3.1.
@@
     $F$ is plateaued with single amplitude if and only if the size of the set in eq. 1 does not depend on $x\in\mathbb{F}_{2}^{n}$, nor on $w\in\mathbb{F}_{2}^{m}$ when $w\neq 0$.
-Theorem 4.3 is very useful for classifying plateaued functions, and in particular, we know that the size of the set
+Hence, for a plateaued function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$, the size of the set
 $\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:F(x)+F(x+a)+F(x+b)+F(x+a+b)=w\right\}$
-is invariant of the choice of $x\in\mathbb{F}_{2}^{n}$ when $w\in\mathbb{F}_{2}^{n}$ is fixed and $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is a plateaued function.
-By replacing $b$ with $x+b$, the size of the above set is the same as
+is invariant of the choice of $x\in\mathbb{F}_{2}^{n}$ when $w\in\mathbb{F}_{2}^{n}$ is fixed. By replacing $b$ with $x+b$, the size of the above set is the same as
 $|\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:F(x)+F(x+a)+F(b)+F(a+b)=w\right\}|$
@@
 $|\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:D_{a}D_{b}F(x)=w\right\}|=\sum_{a\in\mathbb{F}_{2}^{n}}\delta_{F}(a+x,F(a)+F(x)+w).$
-In the particular case that $F$ is also APN, then $\delta_{F}=2\gamma_{F}+2^{n}\delta_{(0,0)}$.
-So when $F$ is APN and $w\neq 0$, we have $\sum_{a\in\mathbb{F}_{2}^{n}}\delta_{F}(a+x,F(a)+F(x)+w)=2\sum_{a\in\mathbb{F}_{2}^{n}}\gamma_{F}(a+x,F(a)+F(x)+w)$, and by Proposition 3.3 we know that $\sum_{a\in\mathbb{F}_{2}^{n}}\gamma_{F}(a+x,F(a)+F(x)+w)=3\operatorname{mult}_{\mathcal{G}_{F}}(x,F(x)+w)$.
+Let us consider the case that $F$ is APN.
+Then $\delta_{F}=2\gamma_{F}+2^{n}\delta_{(0,0)}$.
+So, if $w\neq 0$, then $\sum_{a\in\mathbb{F}_{2}^{n}}\delta_{F}(a+x,F(a)+F(x)+w)=2\sum_{a\in\mathbb{F}_{2}^{n}}\gamma_{F}(a+x,F(a)+F(x)+w)$, and by Proposition 3.3 we know that $\sum_{a\in\mathbb{F}_{2}^{n}}\gamma_{F}(a+x,F(a)+F(x)+w)=3\operatorname{mult}_{\mathcal{G}_{F}}(x,F(x)+w)$.
 Hence
 $|\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:D_{a}D_{b}F(x)=w\right\}|=6\operatorname{mult}_{\mathcal{G}_{F}}(x,F(x)+w)$ (2)
-when $F$ is APN and $w\neq 0$.
-So Theorem 4.3 implies that if $F$ is plateaued APN and $w\neq 0$, then $\operatorname{mult}_{\mathcal{G}_{F}}(x,F(x)+w)$ does not depend on $x\in\mathbb{F}_{2}^{n}$.
-For $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ quadratic, it was shown in [5, Proposition 5] that $\Pi_{F}^{\beta}$ does not depend on $\beta\in\mathbb{F}_{2}^{n}$, and so if $F$ is APN quadratic, the exclude multiplicities $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ as $b$ ranges across $\mathbb{F}_{2}^{n}\setminus\left\{a\right\}$ does not depend on the choice of $a\in\mathbb{F}_{2}^{n}$ by Proposition 3.3.
+for any $w\neq 0$.
+So Theorem 4.3 implies that if $F$ is plateaued and $w\neq 0$, then $\operatorname{mult}_{\mathcal{G}_{F}}(x,F(x)+w)$ does not depend on $x\in\mathbb{F}_{2}^{n}$.
+For $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ quadratic, it was shown in [6, Proposition 5] that $\Pi_{F}^{\beta}$ does not depend on $\beta\in\mathbb{F}_{2}^{n}$, and as a consequence, the exclude multiplicities $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ as $b$ ranges across $\mathbb{F}_{2}^{n}\setminus\left\{F(a)\right\}$ does not depend on the choice of $a\in\mathbb{F}_{2}^{n}$ by Proposition 3.3.
 In the following proposition, we prove this result for all plateaued APN functions.
@@
 For any non-negative integer $k$, let $z_{k}=|\left\{b\in\mathbb{F}_{2}^{n}\setminus\left\{F(0)\right\}:\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=k\right\}|$.
-By Proposition 4.4, we know that for any $a\in\mathbb{F}_{2}^{n}$, there exists a bijection from $\left\{a\right\}\times(\mathbb{F}_{2}^{n}\setminus\left\{F(a)\right\})$ to $\left\{0\right\}\times(\mathbb{F}_{2}^{n}\setminus\left\{F(0)\right\})$ that preserves exclude multiplicity.
+By Proposition 4.4, we know that for any $a\in\mathbb{F}_{2}^{n}$ there exists a bijection from $\left\{a\right\}\times(\mathbb{F}_{2}^{n}\setminus\left\{F(a)\right\})$ to $\left\{0\right\}\times(\mathbb{F}_{2}^{n}\setminus\left\{F(0)\right\})$ that preserves exclude multiplicity.
 Therefore, $m_{k}=2^{n}z_{k}$ for all $k\geq 0$, so $2^{n}|m_{k}$.
 ∎
-From this, we prove that the Brinkmann-Leander-Edel-Pott function $F\colon\mathbb{F}_{2^{6}}\to\mathbb{F}_{2^{6}}$ (see [23], [11, Section 11.5.3], or [4]), given by
+From this, we prove that the Brinkmann-Leander-Edel-Pott function $F\colon\mathbb{F}_{2^{6}}\to\mathbb{F}_{2^{6}}$ (see [24], [14, Section 11.5.3], or [5]), given by
 $F(x)$ $=x^{3}+u^{17}(x^{17}+x^{18}+x^{20}+x^{24})+u^{14}(u^{18}x^{9}+u^{36}x^{18}+u^{9}x^{36}+x^{21}+x^{42})$
@@
 where $u\in\mathbb{F}_{2^{n}}$ is primitive, cannot be CCZ-equivalent to a plateaued function.
-This improves our understanding on this sporadic APN function as the Brinkmann-Leander-Edel-Pott function was only previously known to not be CCZ-equivalent to a quadratic function or monomial function [23].
+This improves our understanding of this sporadic APN function as the Brinkmann-Leander-Edel-Pott function was only previously known to not be CCZ-equivalent to a quadratic function or monomial function [24].
 Corollary 4.6.
@@
 ∎
-Our proof technique for showing that the Brinkmann-Leander-Edel-Pott function is not CCZ-equivalent to a plateaued function also yields a practical test that can be applied to newly discovered APN functions in the future.
+Our proof technique for showing that the Brinkmann-Leander-Edel-Pott function is not CCZ-equivalent to a plateaued function also yields a practical test that can be applied to newly discovered APN functions.
 We can also describe the exclude multiplicities of $\mathcal{G}_{F}$ more directly in terms of the amplitudes of the component functions of $F$.
@@
 $2\sum_{v\in\left\{c\right\}^{\perp}}\lambda_{v}^{2}-\sum_{v\in\mathbb{F}_{2}^{n}}\lambda_{v}^{2}=\sum_{v\in\mathbb{F}_{2}^{n}}(-1)^{v\cdot c}\lambda_{v}^{2}=\sum_{v\in\mathbb{F}_{2}^{n}}\lambda_{v}^{2}-2\sum_{v\notin\left\{c\right\}^{\perp}}\lambda_{v}^{2}.$ (4)
-By using the following proposition, we will be able to rephrase minimizing exclude multiplicity in terms of minimizing the sum of squared amplitudes in any linear hyperplane as the sum $\sum_{v\in\mathbb{F}_{2}^{n}}\lambda_{v}^{2}$ is equal to $2^{n}(3\cdot 2^{n}-2)$ precisely when $F$ is APN.
-Proposition 4.8 ([10, Proposition 9]).
+By using the following proposition, we are able to rephrase minimizing exclude multiplicity in terms of minimizing the sum of squared amplitudes in any linear hyperplane.
+This is because $\sum_{v\in\mathbb{F}_{2}^{n}}\lambda_{v}^{2}=2^{n}(3\cdot 2^{n}-2)$ precisely when $F$ is APN.
+Proposition 4.8 ([13, Proposition 9]).
 Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be an APN plateaued function, and let $\lambda_{v}$ denote the amplitude of $v\cdot F$ for all $v\in\mathbb{F}_{2}^{n}$.
@@
 $\sum_{v\in\mathbb{F}_{2}^{n}}\lambda_{v}^{2}\leq 2^{n}(3\cdot 2^{n}-2).$ (5)
-In particular, if $F$ is APN, then eq. 5 is an equality.
+In particular, if $F$ is APN, then (5) is an equality.
 Corollary 4.9.
@@
 Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued APN function, and let $\lambda_{v}$ denote the amplitude of $v\cdot F$ for all $v\in\mathbb{F}_{2}^{n}$.
-If for a non-negative integer $k$, we have
-$\sum_{v\in\left\{c\right\}^{\perp}}\lambda_{v}^{2}\geq 3\cdot 2^{n}k+2^{n-1}(3\cdot 2^{n}-2)$ (6)
-for all nonzero $c\in\mathbb{F}_{2}^{n}$, then $e_{\min}(\mathcal{G}_{F})\geq k$.
+For any $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, we have
+$\sum_{v\in\left\{b+F(a)\right\}^{\perp}}\lambda_{v}^{2}=3\cdot 2^{n}\operatorname{mult}_{\mathcal{G}_{F}}(a,b)+2^{n-1}(3\cdot 2^{n}-2).$ (6)
 Proof.
-Suppose $k$ is a non-negative integer such that for all nonzero $c\in\mathbb{F}_{2}^{n}$ eq. 6 holds. For any $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, we have by Corollary 4.9 that
-$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ $\geq\frac{1}{3\cdot 2^{n}}\left(3\cdot 2^{n}k+2^{n-1}(3\cdot 2^{n}-2)-2^{n-1}(3\cdot 2^{n}-2)\right)=k.$
-So $e_{\min}(\mathcal{G}_{F})\geq k$. ∎
+Immediate upon rearrangement of the equation from Corollary 4.9. ∎
 Although Corollary 4.10 follows easily from previous statements, it helps shape our approach on finding a lower bound on the minimum exclude multiplicity of the graph of a plateaued APN function.
@@
 We use the following corollary which follows from a classical result of [4].
-Corollary 4.11 ([34, Corollary 1]).
+Corollary 4.11 ([35, Corollary 1]).
 A set $S\subseteq(\mathbb{F}_{2}^{n}\setminus\left\{0\right\})$ that intersects every (n+1k)(n+1-k)-dimensional subspace of $\mathbb{F}_{2}^{n}$ has at least $2^{k}-1$ elements with equality if and only if $S\cup\left\{0\right\}$ is a $k$-dimensional subspace of $\mathbb{F}_{2}^{n}$.
@@
 Proof.
-In [34, Proof of Theorem 3], it was shown that there cannot exist a $(\frac{n}{2}+1)$-dimensional linear subspace $T$ of $\mathbb{F}_{2}^{n}$ such that $T\setminus\left\{0\right\}$ is contained in $\mathcal{B}({F})$} (otherwise we could easily construct a function $\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n/2+1}$ whose component functions are all bent, but this is impossible by Nyberg’s bound).
+In [35, Proof of Theorem 3], it was shown that there cannot exist a $(\frac{n}{2}+1)$-dimensional linear subspace $T$ of $\mathbb{F}_{2}^{n}$ such that $T\setminus\left\{0\right\}$ is contained in $\mathcal{B}({F})$} (otherwise we could easily construct a function $\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n/2+1}$ whose component functions are all bent, but this is impossible by Nyberg’s bound).
 Hence, $S\cap{\mathcal{B}({F})}$ cannot contain all nonzero elements of a linear subspace of dimension $\frac{n}{2}+1$ because it is a subset of $\mathcal{B}({F})$}.
@@
 with equality if and only if $S\setminus{\mathcal{B}({F})}$ is a $(d-\frac{n}{2})$-dimensional subspace of $S$. ∎
-From this, we are able to derive lower bounds on the exclude multiplicities of the graphs of APN functions (for $n$ even).
+From this, we are able to derive lower bounds on the exclude multiplicities of the graphs of plateaued APN functions (for $n$ even).
 Theorem 4.13.
-Assume $n\geq 4$ is even.
+Assume $n$ is even.
 Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued APN function.
@@
 Proof.
-Let $c=F(a)+b$ for $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$. Let $N=|\left\{c\right\}^{\perp}\setminus({\mathcal{B}({F})}\cup\left\{0\right\})|$. By applying Corollary 4.9, we have
+Let $c=F(a)+b$ for $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$.
+Let $N=|\left\{c\right\}^{\perp}\setminus({\mathcal{B}({F})}\cup\left\{0\right\})|$.
+Recall that if $v\in{\mathcal{B}({F})}$, then $\lambda_{v}=2^{\frac{n}{2}}$, and otherwise, $\lambda_{v}\geq 2^{\frac{n}{2}+1}$.
+So
 $\sum_{v\in\left\{c\right\}^{\perp}}\lambda_{v}^{2}$ $=2^{2n}+(2^{n-1}-1-N)2^{n}+\sum_{v\in\left\{c\right\}^{\perp}\setminus({\mathcal{B}({F})}\cup\left\{0\right\})}\lambda_{v}^{2}$
@@
 By Lemma 4.12, we have $N\geq 2^{\frac{n}{2}-1}-1$, so $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq e_{\min}(\mathcal{G}_{F})\geq N\geq 2^{\frac{n}{2}-1}-1$ by Corollary 4.10. ∎
-Our lower bound in Theorem 4.13 is a dramatic increase over the previously known $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq 1$ when $F$ is plateaued APN.
+Our lower bound in Theorem 4.13 is a dramatic increase over the previously known $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq 1$ (for $n\geq 3$) when $F$ is plateaued APN.
 It would be interesting to further improve this lower bound or find an example of a family of plateaued APN functions for which it is tight.
@@
 However, when $n$ is even, the exclude multiplicities of $\mathcal{G}_{F}$ are much more unclear.
-Computational results seem to suggest that if $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is a plateaued APN function, then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd for all $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, and we conjecture in 4.15 that this is true.
 In the following proposition, we describe the parity of the exclude multiplicities of $\mathcal{G}_{F}$ for $F$ plateaued APN in terms of the Fourier transform of the indicator function $1_{\mathcal{B}({F})}\colon\mathbb{F}_{2}^{n}\to\left\{0,1\right\}$ of $\mathcal{B}({F})$}.
+This then tells us that the exclude multiplicities of the graph of any plateaued function are always odd (for $n\geq 3$ since $\mathcal{G}_{F}$ is not a maximal Sidon set if $n\leq 2$).
 Proposition 4.14.
-Assume $n$ is even.
-Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be plateaued APN function and let $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$.
-Then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd if and only if $\widehat{1_{\mathcal{B}({F})}}(F(a)+b)\equiv 2\mod 4$.
-Equivalently, $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd if and only if $|{\mathcal{B}({F})}\cap\left\{F(a)+b\right\}^{\perp}|$ is even.
+Assume $n\geq 3$.
+Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be plateaued APN function, and let $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$.
+Then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd.
 Proof.
+If $n$ is odd, then $F$ is AB and $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd.
+So, assume $n$ is even.
+We claim that $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd if and only if $\widehat{1_{\mathcal{B}({F})}}(F(a)+b)\equiv 2\mod 4$.
 For all $v\in\mathbb{F}_{2}^{n}$, let $\lambda_{v}$ denote the amplitude of $v\cdot F$, and let $m_{v}$ be the non-negative integer such that $\lambda_{v}^{2}=2^{n+m_{v}}$.
@@
 However, we already know that $\widehat{1_{\mathcal{B}({F})}}(c)+\widehat{1_{\mathbb{F}_{2}^{n}\setminus{\mathcal{B}({F})}}}(c)=0$ as $c\neq 0$.
-So $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd if and only if $\widehat{1_{\mathcal{B}({F})}}(c)\equiv 2\mod 4$, and this proves the first part of the statement.
-For $i\in\left\{0,1\right\}$, let $B_{i}=|\left\{v\in{\mathcal{B}({F})}:v\cdot c=i\right\}|$.
+So $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd if and only if $\widehat{1_{\mathcal{B}({F})}}(c)\equiv 2\mod 4$, and this proves our claim.
+Now, for $i\in\left\{0,1\right\}$, let $B_{i}=|\left\{v\in{\mathcal{B}({F})}:v\cdot c=i\right\}|$.
 Then $\widehat{1_{\mathcal{B}({F})}}(c)=|B_{0}|-|B_{1}|=|{\mathcal{B}({F})}|-2|B_{1}|$.
-It was shown in [30, Proposition 6.5] that the number of bent components of any plateaued function over an even dimension is $2\mod 4$, and so $|{\mathcal{B}({F})}|\equiv 2\mod 4$.
+It was shown in [31, Proposition 6.5] that the number of bent components of any plateaued function over an even dimension is $2\mod 4$, and so $|{\mathcal{B}({F})}|\equiv 2\mod 4$.
 Therefore, $\widehat{1_{\mathcal{B}({F})}}(c)\equiv 2-2|B_{1}|\mod 4$.
 So $\widehat{1_{\mathcal{B}({F})}}(c)\equiv 2\mod 4$ if and only if $|B_{1}|$ is even.
-Since the parity of $|B_{0}|$ and $|B_{1}|$ is the same, the second part of the result follows.
+Since the parity of $|B_{0}|$ and $|B_{1}|$ is the same, it follows that $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd if and only if $|{\mathcal{B}({F})}\cap\left\{F(a)+b\right\}^{\perp}|$ is even.
+Recently, it was shown in [1, Proposition 3.1] that for any plateaued APN function with $n$ even, the intersection of $\mathbb{F}_{2}^{n}\setminus({\mathcal{B}({F})}\cup\left\{0\right\})$ with any subspace of dimension at least $\frac{n}{2}+1$ is odd.
+In particular, this implies that $|{\mathcal{B}({F})}\cap\left\{F(a)+b\right\}^{\perp}|$ is even.
 ∎
-Conjecture 4.15.
-Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be an APN plateaued function.
-Then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd for all $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$.
-In Section 4.4, we prove the above conjecture in the affirmative in the case of $F$ being quadratic APN, but we were unable to find a proof for all plateaued APN functions.
-We now express the exclude multiplicities of the graph of an APN function in terms of $\widehat{W_{F}^{4}}$.
+In Section 4.4, we provide another proof of the above proposition for quadratic APN functions via the ortho-derivative.
+We now express the exclude multiplicities of the graph of a plateaued APN function in terms of $\widehat{W_{F}^{4}}$.
 Proposition 4.15.
@@
 $\frac{1}{2^{2n}}\widehat{W_{F}^{4}}(0,c)$ $=|\left\{(x,y,z,w)\in(\mathbb{F}_{2}^{n})^{4}:(x+y+z+w,F(x)+F(y)+F(z)+F(w))=(0,c)\right\}|$
-$=|\left\{(x,y,z)\in(\mathbb{F}_{2}^{n})^{3}:F(x)+F(y)+F(z)+F(x+y+z)=(0,c)\right\}|$
+$=|\left\{(x,y,z)\in(\mathbb{F}_{2}^{n})^{3}:F(x)+F(y)+F(z)+F(x+y+z)=c\right\}|$
 $=|\left\{(x,a,b)\in(\mathbb{F}_{2}^{n})^{3}:F(x)+F(x+a)+F(x+b)+F(x+a+b)=c\right\}|$
@@
 In Section 7, we will see why this is relevant and how it is connected to the existence of nontrivial linear structures of $\gamma_{F}$.
-In [13, Section 6], it was shown that for any APN function $F$, we have $W_{F}^{2}=2^{2n}\delta_{(0,0)}-W_{\gamma_{F}}+2^{n}$, (where $\delta_{(0,0)}$ takes value $1$ at $(0,0)$ and 0 elsewhere).
+In [15, Section 6], it was shown that for any APN function $F$, we have $W_{F}^{2}=2^{2n}\delta_{(0,0)}-W_{\gamma_{F}}+2^{n}$, (where $\delta_{(0,0)}$ takes value $1$ at $(0,0)$ and 0 elsewhere).
 So
@@
 We say a function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is 33-to-11 if every point in the image of $F$ has a preimage of size $3$ except a single point.
-Note that all 3-to-1 plateaued functions are necessarily APN (see [29, Theorem 12]).
+Note that all 3-to-1 plateaued functions are necessarily APN (see [30, Corollary 8]).
 A subset $D\subseteq G$ of an additive group $G$ is a partial difference set with parameters $(v,k,\lambda,\mu)$ if every non-identity element of $D$ (resp.
@@
 The following theorem was originally only proved for crooked 33-to-11 functions (in short, those functions such that $\operatorname{Im}D_{a}F$ is an affine hyperplane for all $a\neq 0$), but its proof only relied on the fact that all crooked functions are plateaued, so we write it for all 33-to-11 plateaued functions.
-Theorem 4.18 ([7, Theorem 4]).
+Theorem 4.17 ([8, Theorem 4]).
 Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued such that $F(0)=0$ and $F$ is 3-to-1 on $\mathbb{F}_{2}^{n}\setminus\left\{0\right\}$.
 Then $\operatorname{Im}F\setminus\left\{0\right\}$ is a partial difference set, with parameters $(2^{n},\frac{2^{n}-1}{3},\frac{2}{3}(\alpha(n)-1),\frac{2}{3}\beta(n))$, where $\alpha(n)=\frac{2^{n}+(-2)^{\frac{n}{2}+1}-2}{6}$ and $\beta(n)=\frac{2^{n}+(-2)^{\frac{n}{2}}-2}{6}$.
-Moreover, [7, Corollary 4] describes the multiplicities of the nonzero elements of the multiset $\left\{F(x)+F(y)+F(x+y):x,y\in\mathbb{F}_{2}^{n}\right\}$ when $F$ is quadratic 33-to-11, and indeed, this can be easily generalized to all plateaued 3-to-1 APN functions.
+Moreover, [8, Corollary 4] describes the multiplicities of the nonzero elements of the multiset $\left\{F(x)+F(y)+F(x+y):x,y\in\mathbb{F}_{2}^{n}\right\}$ when $F$ is quadratic 33-to-11, and indeed, this can be easily generalized to all plateaued 3-to-1 APN functions.
 Corollary 4.18.
@@
 Then $\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=\alpha(n)$ if $b\in\operatorname{Im}F$ and $\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=\beta(n)$ if $b\notin\operatorname{Im}F$.
-These two results above (in terms of quadratic 3-to-1 functions) and [5, Proposition 5] (which states that $\Pi_{F}^{\beta}$ does not depend on $\beta\in\mathbb{F}_{2}^{n}$ when $F$ is quadratic) were used in [7] to determine the exact exclude multiplicities of $\mathcal{G}_{F}$ and their frequencies in the case of $F$ being a quadratic 33-to-11 function.
+These two results above (in terms of quadratic 3-to-1 functions) and [6, Proposition 5] (which states that $\Pi_{F}^{\beta}$ does not depend on $\beta\in\mathbb{F}_{2}^{n}$ when $F$ is quadratic) were used in [8] to determine the exact exclude multiplicities of $\mathcal{G}_{F}$ and their frequencies in the case of $F$ being a quadratic 33-to-11 function.
 This can be done because any plateaued function $F$ with all component functions unbalanced satisfies
 $|\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:D_{a}D_{b}F(x)=w\right\}|=|\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:F(a)+F(b)=w\right\}|$ (8)
-for all $x,w\in\mathbb{F}_{2}^{n}$ (see [10, Theorem 2]). Therefore, if $F$ is plateaued APN with all unbalanced components and $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, then
+for all $x,w\in\mathbb{F}_{2}^{n}$ (see [13, Theorem 2]). Therefore, if $F$ is plateaued APN with all unbalanced components and $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$, then
 $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\frac{1}{6}|\left\{(u,v)\in(\mathbb{F}_{2}^{n})^{2}:F(u)+F(v)=F(a)+b\right\}|$
@@
 The result then immediately follows by applying Proposition 4.4 and Corollary 4.18 (cf.
-[7]) and using the same proof as [7, Corollary 4].
+[8]) and using the same proof as [8, Corollary 4].
 ∎
@@
 Therefore if $F$ is a quadratic APN function with $F(0)=0$ and $b\neq 0$, then
-$\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=\frac{1}{6\cdot 2^{2n}}\widehat{W_{F}^{3}}(0,b)=\frac{1}{6\cdot 2^{3n}}\widehat{W_{F}^{4}}(0,b)=\frac{1}{3}(2^{n}-1-\operatorname{wt}(b\cdot\pi_{F})).$ (9)
+$\operatorname{mult}_{\mathcal{G}_{F}}(0,b)=\frac{1}{6\cdot 2^{2n}}\widehat{W_{F}^{3}}(0,b)=\frac{1}{6\cdot 2^{3n}}\widehat{W_{F}^{4}}(0,b)=\frac{1}{3}(2^{n}-1-\operatorname{wt}(b\cdot\pi_{F}))$ (9)
+with the second equality following from Proposition 4.15.
 Remark 4.22.
@@
 Proposition 4.23.
-Assume $n\geq 4$ is even.
+Assume $n$ is even.
 Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a quadratic APN function with $F(0)=0$.
@@
 For any function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$, it is well-known that $F$ has algebraic degree $n$ if and only if $\sum_{x\in\mathbb{F}_{2}^{n}}F(x)\neq 0$.
-Therefore, if $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is quadratic APN, then $\sum_{x\in\mathbb{F}_{2}^{n}}\pi_{F}(x)=0$, implying that all component functions of $\pi_{F}$ have even Hamming weight as for any $v\neq 0$, we have $\operatorname{wt}(v\cdot\pi_{F})\equiv\sum_{x\in\mathbb{F}_{2}^{n}}v\cdot\pi_{F}(x)\pmod{2}=v\cdot\sum_{x\in\mathbb{F}_{2}^{n}}\pi_{F}(x)=0.$ This allows us to prove that the exclude multiplicities of the graph of an APN quadratic function are always odd.
+Therefore, if $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ is quadratic APN, then $\sum_{x\in\mathbb{F}_{2}^{n}}\pi_{F}(x)=0$, implying that all component functions of $\pi_{F}$ have even Hamming weight as for any $v\neq 0$, we have $\operatorname{wt}(v\cdot\pi_{F})\equiv\sum_{x\in\mathbb{F}_{2}^{n}}v\cdot\pi_{F}(x)\pmod{2}=v\cdot\sum_{x\in\mathbb{F}_{2}^{n}}\pi_{F}(x)=0.$ This allows us to prove that the exclude multiplicities of the graph of an quadratic APN function are always odd in a different manner than Proposition 4.14.
 Proposition 4.24.
@@
 Then $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)$ is odd for all $(a,b)\in(\mathbb{F}_{2}^{n})^{2}\setminus\mathcal{G}_{F}$.
-Hence, $|\mathcal{B}(F)\cap H|$ is even for any affine hyperplane $H$ of $\mathbb{F}_{2}^{n}$.
+Equivalently, $|\mathcal{B}(F)\cap H|$ is even for any affine hyperplane $H$ of $\mathbb{F}_{2}^{n}$.
 Proof.
@@
 Thus, by applying Proposition 4.4, we know that any exclude point of $\mathcal{G}_{F}$ has odd multiplicity.
-The parity of $|\mathcal{B}(F)\cap H|$ must then be even for any affine hyperplane $H$ of $\mathbb{F}_{2}^{n}$, see Proposition 4.14.
 ∎
@@
 As a corollary of Proposition 4.24, every quadratic APN function has a maximal Sidon set as its graph, and as previously mentioned, this is already known since all quadratic functions are plateaued.
-Nonetheless, it is interesting that Proposition 4.25 now provides a fourth method (the other three being the work of [6], the usage of the D-property in [12], and Theorem 4.13) for proving no quadratic APN function has a non-maximal Sidon set as its graph.
+Nonetheless, it is interesting that Proposition 4.24 now provides a fourth method for proving no quadratic APN function (for $n\geq 3$) has a non-maximal Sidon set as its graph (the other three methods being the work of [7], the usage of the D-property in [16], and Theorem 4.13).
 5 Power functions
@@
 ∎
-If $F(x)=x^{d}$ is a APN and $n$ is odd, we can determine the exact exclude multiplicities of $\mathcal{G}_{F}$ contained in $\left\{0\right\}\times\mathbb{F}_{2^{n}}^{\ast}$, but first, we describe the sum $\sum_{x\in\mathbb{F}_{2^{n}}}\gamma_{F}(x+a,x^{d}+b)$ in the following way to make our computation straightforward.
+If $F(x)=x^{d}$ is APN and $n$ is odd, we can determine the exact exclude multiplicities of $\mathcal{G}_{F}$ contained in $\left\{0\right\}\times\mathbb{F}_{2^{n}}^{\ast}$, but first we describe the sum $\sum_{x\in\mathbb{F}_{2^{n}}}\gamma_{F}(x+a,x^{d}+b)$ in the following way to make our computation straightforward.
 Proposition 5.2.
@@
 $\operatorname{mult}_{\mathcal{G}_{F}}(0,a)=\frac{|\Phi^{-1}(\operatorname{Im}D_{1}F)|}{3}=\frac{2^{n-1}-1}{3}.$
-In [27, Conjecture 21], Kaleyski conjectured that if $F$ is an APN power function, the two values $\sum_{y\in\mathbb{F}_{2^{n}}}\delta_{F}(y,F(y)+1)$ and $\sum_{y\in\mathbb{F}_{2^{n}}}\delta_{F}(y+1,F(y))$ are equal to $\sum_{y\in\mathbb{F}_{2^{n}}}\delta_{x^{3}}(y,y^{3}+1)$.
+In [28, Conjecture 21], Kaleyski conjectured that if $F$ is an APN power function, the two values $\sum_{y\in\mathbb{F}_{2^{n}}}\delta_{F}(y,F(y)+1)$ and $\sum_{y\in\mathbb{F}_{2^{n}}}\delta_{F}(y+1,F(y))$ are equal to $\sum_{y\in\mathbb{F}_{2^{n}}}\delta_{x^{3}}(y,y^{3}+1)$.
 By Proposition 3.3, this is equivalent to $\operatorname{mult}_{\mathcal{G}_{F}}(0,1)=\operatorname{mult}_{\mathcal{G}_{F}}(1,0)=\operatorname{mult}_{\mathcal{G}_{x^{3}}}(0,1)$.
@@
 Kloosterman sums can also be used to describe the number of rational points on ordinary elliptic curves over $\mathbb{F}_{2^{n}}$.
-First, recall that an algebraic plane curve $\mathcal{C}$ over $\mathbb{F}_{2^{n}}$ is defined by the equation $p(x,y)=0$ for some irreducible polynomial $p(x,y)$ over $\mathbb{F}_{2^{n}}$.
+Recall that an algebraic plane curve $\mathcal{C}$ over $\mathbb{F}_{2^{n}}$ is defined by the equation $p(x,y)=0$ for some irreducible polynomial $p(x,y)$ over $\mathbb{F}_{2^{n}}$.
 If $\widetilde{p}(x,y,z)$ is the associated homogenized polynomial to $p(x,y)$ we denote the solution set to $\widetilde{p}(x,y,z)=0$ as $\widetilde{C}$.
@@
 Then for all $(a,b)\in\mathbb{F}_{2^{n}}^{2}\setminus\mathcal{G}_{F}$, we have
-$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq\left\lceil\frac{2^{n}-5-\max_{a\in\mathbb{F}_{2^{n}}^{\ast}}|K_{n}(a)-1|}{6}\right\rceil$ (10)
+$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq\left\lceil\frac{2^{n}-5-\max_{a\in\mathbb{F}_{2^{n}}^{\ast}}|K_{n}(a)-1|}{6}\right\rceil.$ (10)
 Proof.
@@
 $q(x,y)=bx^{2}y+bxy^{2}+x^{2}+y^{2}+bxy+xy+x+y.$ (12)
-Let $\mathcal{E}=\{(x,y)\in\mathbb{F}_{2^{n}}^{2}:q(x,y)=0\}$, and let $G=\left\{(x,y)\in E:0\notin\left\{x,y,x+y+1\right\}\right\}$.
+Let $\mathcal{E}=\{(x,y)\in\mathbb{F}_{2^{n}}^{2}:q(x,y)=0\}$, and let $G=\left\{(x,y)\in\mathcal{E}:0\notin\left\{x,y,x+y+1\right\}\right\}$.
 Note that for all $(x,y)\in\mathbb{F}_{2^{n}}^{2}$ such that $x,y,x+y+1\neq 0$, we have $p(x,y)=0$ if and only if $q(x,y)=0.$ Hence $|D|=|G|$.
@@
 Note that the projective points of $\widetilde{\mathcal{E}}$ include three additional points, corresponding to solutions of the form $(x:y:0)$ that satisfy $\widetilde{q}(x,y,0)=bx^{2}y+bxy^{2}=bxy(x+y)=0$.
-There are three such cases: $x=0,y=0$ and $x+y=0,$ corresponding to the points $(1:0:0),(0:1:0),(1:1:0)\in\widetilde{E}$ respectively.
+There are three such cases: $x=0,y=0$ and $x+y=0,$ corresponding to the points $(1:0:0),(0:1:0),(1:1:0)\in\widetilde{\mathcal{E}}$ respectively.
 So $\widetilde{\mathcal{E}}=\#\mathcal{E}+3$.
@@
 However, not very much is known about whether or not $\gamma_{F}$ can have 0-valued linear structures.
-In [13], there were two classes of APN functions that were shown that do not admit linear structures: the first if when $F$ is an APN power function, and the second is when $F$ is such that there exists some $x\in\mathbb{F}_{2}^{n}$ with the property that $\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:D_{a}D_{b}F(x)=w\right\}$ does not depend on $w\in\mathbb{F}_{2}^{n}\setminus\left\{0\right\}$.
-We will prove that if $F$ is a plateaued 3-to-1 function ($n$ even), then $\gamma_{F}$ does not admit nontrivial linear structures, forming the third known class of APN functions such that $\gamma_{F}$ has no nontrivial linear structures.
+In [15], two classes of APN functions were given such that $\gamma_{F}$ does not admit non-trivial linear structures.
+The first is when $F$ is an APN power function, and the second is when $F$ is such that there exists some $x\in\mathbb{F}_{2}^{n}$ with the property that the size of $\left\{(a,b)\in(\mathbb{F}_{2}^{n})^{2}:D_{a}D_{b}F(x)=w\right\}$ does not depend on $w\in\mathbb{F}_{2}^{n}\setminus\left\{0\right\}$.
+We prove that if $F$ is a plateaued APN function with $n$ even, then $\gamma_{F}$ does not admit nontrivial linear structures, forming the third known class of APN functions such that $\gamma_{F}$ has no nontrivial linear structures (excluding AB functions as $\gamma_{F}$ is bent in this case and therefore has balanced nonzero derivatives).
 Let us rephrase the problem of $\gamma_{F}$ having linear structures in terms of exclude multiplicity when $F$ is a plateaued APN function.
-Proposition 7.1.
+Lemma 7.1.
 Assume $n\geq 4$ is even.
@@
 Therefore $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)=\frac{2^{n}-1}{3}$ if and only if $\operatorname{wt}(D_{(0,b+F(a))}\gamma_{F})=0$. ∎
-Proposition 7.1 shows that obtaining an upper bound on the exclude multiplicities of the graph of a plateaued APN function also poses as a difficult problem.
-However, recall that in Section 4.3, we determined the exact exclude multiplicities that occur when $F$ is a plateaued 3-to-1 function.
-The following result is then an immediate corollary.
+The following result is then an immediate corollary of the fact that plateaued APN functions cannot have an algebraic degree equal to $n$, see [7].
 Proposition 7.2.
-Assume $n\geq 4$ and is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued 3-to-1 function. Then $\gamma_{F}$ has no nontrivial linear structures.
+Assume $n\geq 4$ is even. Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued APN function. Then $\gamma_{F}$ has no nontrivial linear structures.
 Proof.
-By Theorem 4.20, any exclude point of $\mathcal{G}_{F}$ has multiplicity $\frac{2^{n}+(-2)^{\frac{n}{2}+1}-2}{6}$ or $\frac{2^{n}+(-2)^{\frac{n}{2}}-2}{6}$, and so any exclude point of $\mathcal{G}_{F}$ has multiplicity strictly less than $\frac{2^{n}-1}{3}$.
-The result then follows from Proposition 7.1.
+Assume that $F$ is plateaued, and by way of contradiction assume that $(0,F(0)+b)$ has an exclude multiplicity of $\frac{2^{n}-1}{3}$ for some $b\neq 0$.
+Then $\mathbb{F}_{2}^{n}\setminus\{0\}$ decomposes into $\frac{2^{n}-1}{3}$ disjoint triples $\{x,y,z\}$ with
+$(x+y+z,F(x)+F(y)+F(z))=(0,b+F(0)).$
+Summing over all triples gives $\sum_{x\in\mathbb{F}_{2}^{n}\setminus\left\{0\right\}}(x,F(x))=(0,b+F).$ Moreover, since $F$ has algebraic degree strictly less than $n$, we know $\sum_{x\in\mathbb{F}_{2}^{n}}F(x)=0$, implying $\sum_{x\in\mathbb{F}_{2}^{n}\setminus\left\{0\right\}}(x,F(x))=(0,F(0))$.
+Therefore, $b=0$, a contradiction.
+Thus, no exclude point of $\mathcal{G}_{F}$ can have multiplicity equal to $\frac{2^{n}-1}{3}$ by Proposition 4.4, and the result follows from Lemma 7.1.
 ∎
-As shown above, there is an equivalence of $\gamma_{F}$ having a nontrivial linear structure and $\mathcal{G}_{F}$ having an exclude point of multiplicity $\frac{2^{n}-1}{3}$ when $F$ is plateaued APN.
-In the case of $F$ also being quadratic, we can then characterize the existence of a nontrivial linear structure of $\gamma_{F}$ in terms of the components of the ortho-derivative of $F$.
+As shown above, there is an equivalence of $\gamma_{F}$ having a nontrivial linear structure and $\mathcal{G}_{F}$ having an exclude point of multiplicity $\frac{2^{n}-1}{3}$ when $F$ is plateaued APN, and we showed that this is an impossibility.
+In the case of $F$ also being quadratic APN, we can then demonstrate that the ortho-derivative of $F$ cannot have trivial component functions.
 Corollary 7.3.
 Suppose $n\geq 4$ is even, and let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a quadratic APN function.
-Then $\gamma_{F}$ has a nontrivial linear structure if and only if the ortho-derivative of $F$ has a component function of weight 0.
-Equivalently, $\gamma_{F}$ has a nontrivial linear structure if and only if the image of $\pi_{F}$ is contained in a linear hyperplane.
+Then $\pi_{F}$ has no components of weight 0, or equivalently, the image of $\pi_{F}$ is not contained in a linear hyperplane.
 Proof.
-Apply Proposition 7.1 to eq. 9. ∎
-In particular, the class of quadratic APN functions whose ortho-derivatives have nonzero nonlinearity is a subclass of those APN functions such that $\gamma_{F}$ has no nontrivial linear structures.
-This immediately follows by applying the inequality $\operatorname{wt}(b\cdot F)\geq\operatorname{\mathcal{NL}}(\pi_{F})$ for $b\neq 0$ to eq.
-9.
-In particular, we have the inequality
-$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\leq\frac{2^{n}-1-\operatorname{\mathcal{NL}}(\pi_{F})}{3}.$ (13)
+Apply Lemma 7.1 and Proposition 7.2 to eq. 9. ∎
+We now apply the inequality $\operatorname{wt}(b\cdot F)\geq\operatorname{\mathcal{NL}}(\pi_{F})$ for $b\neq 0$ to eq. 9. In particular, we have
+$\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\leq\frac{2^{n}-1-\operatorname{\mathcal{NL}}(\pi_{F})}{3}$ (13)
 when $F$ is quadratic APN and $(a,b)\notin\mathcal{G}_{F}$.
@@
 Suppose $n\geq 4$ is even.
-Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a quadratic APN function with $F(0)=0$.
+Let $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ be a plateaued APN function with $F(0)=0$.
 Then for any $b\neq 0$, the weight of $D_{(0,b)}\gamma_{F}$ is divisible by $6\cdot 2^{n}$ and satisfies
@@
 So $\operatorname{wt}(D_{(0,b)}\gamma_{F})=2^{n}(2^{n}-6k-4)$.
-Since $2^{n}-4$ is divisible by $6$ as $n$ is even, we have $\operatorname{wt}(D_{(0,b)}\gamma_{F})\equiv 0\mod 6\cdot 2^{n}$.
-By Theorem 4.13, we have $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq 2^{\frac{n}{2}-1}-1$ for all $(a,b)\notin\mathcal{G}_{F}$ as $F$ is quadratic and therefore plateaued.
-Moreover, we have $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\leq\frac{2^{n}-1}{3}$ by [19, Corollary 3.4].
+Since $n$ is even, $2^{n}-4$ is divisible by $6$, and so we have $\operatorname{wt}(D_{(0,b)}\gamma_{F})\equiv 0\mod 6\cdot 2^{n}$.
+By Theorem 4.13, we have $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\geq 2^{\frac{n}{2}-1}-1$ for all $(a,b)\notin\mathcal{G}_{F}$.
+Moreover, we have $\operatorname{mult}_{\mathcal{G}_{F}}(a,b)\leq\frac{2^{n}-1}{3}$ by [20, Corollary 3.4].
 So, for $b\neq 0$, we have $2^{\frac{n}{2}-1}-1\leq\operatorname{mult}_{\mathcal{G}_{F}}(0,b)\leq\frac{2^{n}-1}{3}$, and the result follows.
@@
 Open Problem 8.3.
-For $n$ even, prove or disprove the existence of a plateaued (or quadratic) APN functions whose graph has an exclude point of multiplicity $\frac{2^{n}-1}{3}$ (the maximum that is theoretically possible).
-Recall that Proposition 7.1 characterizes this problem in terms of the existence of a nontrivial linear structure of $\gamma_{F}$.
+Study the nonlinearity of the ortho-derivative $\pi_{F}$ of a quadratic APN function $F$. Is it possible for $\pi_{F}$ to have zero nonlinearity?
 Open Problem 8.4.
-Study the nonlinearity of the ortho-derivative $\pi_{F}$ of a quadratic APN function $F$. Is it possible for $\pi_{F}$ to have zero nonlinearity?
-Open Problem 8.5.
-For $n$ even, find an example of a non-quadratic APN function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ whose graph has an exclude point of even multiplicity, or prove such a function does not exist.
+For $n$ even, find an example of a non-plateaued APN function $F\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n}$ whose graph has an exclude point of even multiplicity, or prove such a function does not exist.
 Note that if such a function does not exist, then 2.1 follows for all even $n$.
-It may be easier to first prove or disprove 4.15.
-Acknowledgment
+Declarations
+Ethics approval and consent to participate:
+Not applicable.
+Consent for publication:
+Not applicable.
+Availability of data and materials:
+This study does not generate or analyze datasets.
+Competing interests:
+The authors declare they have no competing interests.
+Funding:
+This research received no external funding.
+Authors’ contributions:
+All authors contributed equally to this work.
+Acknowledgments:
 We thank Claude Carlet and Larry Rolen for useful discussions and helpful feedback that improved earlier drafts of this paper.